### v1

# A Sampling-based Framework for Hypothesis Testing on Large Attributed Graphs [Scalable Data Science]

Abstract.
Hypothesis testing is a statistical method used to draw conclusions about populations from sample data, typically represented in tables. With the prevalence of graph representations in real-life applications, hypothesis testing in graphs is gaining importance. In this work, we formalize node, edge, and path hypotheses in attributed graphs. We develop a sampling-based hypothesis testing framework, which can accommodate existing hypothesis-agnostic graph sampling methods. To achieve accurate and efficient sampling, we then propose a Path-Hypothesis-Aware SamplEr, PHASE, an $m$-dimensional random walk that accounts for the paths specified in a hypothesis. We further optimize its time efficiency and propose $\text{PHASE}_{\text{opt}}$. Experiments on real datasets demonstrate the ability of our framework to leverage common graph sampling methods for hypothesis testing, and the superiority of hypothesis-aware sampling in terms of accuracy and time efficiency.
1. Introduction
Hypothesis testing finds widespread application in various domains such as marketing, healthcare, and social science (Newey and McFadden, 1994). Hypotheses are usually tested on representative samples since it is impractical or even impossible to extract data from entire populatations due to size, accessibility or cost. For instance, snowball sampling has been proven effective in accessing a hidden population like drug abusers through a chain-referral mechanism (Waters, 2015; Goodman, 1961). In this paper, we study hypothesis testing on graphs.
Graphs can be used to represent many real-world applications such as social, bibliographic and transportation networks where entities are nodes and edges reflect relationships between entities. Specifically, we work on Attributed Graphs that are graphs with multiple types of nodes and edges, both of which are labeled with attributes containing valuable information for graph analytics (Croft et al., 2011; Omidvar-Tehrani et al., 2019). In a bibliographic network like DBLP (see Figure 1a and Figure 1b for an example and an attributed path of length one, respectively), hypotheses expressed on nodes, edges, and paths unveil interesting aspects of collaboration behavior and research trends based on citation evolution over time. Specifically, a node hypothesis can be “the average citation of conference papers is greater than average”, and a path hypothesis can be “the minimum number of citations of conference papers written by Microsoft researchers is greater than average”. Similarly, on social platforms like Yelp, hypothesis testing can be used to check how different users review a product and study its adoption.
Figure 1. DBLP network schema and paths
Objectives and Challenges. Given an input attributed graph $\mathcal{G}$, our aim is to verify a hypothesis $H$ and return true or false. We hence seek to achieve two objectives: O1 - hypothesis testing should be as accurate as possible; O2 - execution time should be as small as possible. To the best of our knowledge, there is no existing literature addressing hypotheses on attributed graphs. A straightforward approach is to obtain a sampled graph $\mathcal{S}$ from $\mathcal{G}$ using hypothesis-agnostic graph sampling methods (Maiya and Berger-Wolf, 2010; Li et al., 2019; Adamic et al., 2001). However, the challenge is that depending on the sampling method and on the budget $B$, i.e., the size of $\mathcal{S}$, the hypothesis estimator, which computes aggregated values of nodes, edges, or paths specified in the hypothesis may return different values on $\mathcal{S}$ and $\mathcal{G}$. This raises the accuracy challenge behind O1. Additionally, as $B$ increases, the accuracy of the resulting hypotheses increases. However, the speed at which the accuracy converges highly depends on the sampling. This raises the efficiency challenge behind O2.
Contributions. We classify hypotheses on attributed graphs into three types: node, edge, and path hypotheses. We develop a sampling-based hypothesis testing framework, which accommodates common hypothesis-agnostic sampling methods, such as random node sampler (Stumpf et al., 2005), simple random walk (Gjoka et al., 2010), and non-backtracking random walk (Lee et al., 2012).
The lack of awareness of the input hypothesis in existing sampling methods may slow down the convergence speed of accuracy. Indeed, the sampled graph $\mathcal{S}$ should preserve as many relevant paths according to $H$ from $\mathcal{G}$ as possible for a path hypothesis. Therefore, to address objectives O1 and O2, we design PHASE, a Path-Hypothesis-Aware SamplEr. PHASE is aware of $H$ and preserves its nodes, edges, or paths. Consequently, the resulting $\mathcal{S}$ is more likely to provide relevant information for testing the hypothesis. PHASE employs $m\geq 1$ dependent random walks with two weight functions to ensure path-hypothesis-awareness. One function biases the seed selection toward the first node in the path hypothesis. The other biases the transition probability towards the nodes specified in the hypothesis. We further propose $\text{PHASE}_{\text{opt}}$ to improve the time efficiency of PHASE. $\text{PHASE}_{\text{opt}}$ reduces computation overhead through random sampling of neighbors. It also adopt a non-backtracking approach to prioritize the selection of relevant nodes, edges, and paths. We demonstrate both theoretically and empirically that the hypothesis estimator on $\mathcal{S}$ converges to its corresponding value in $\mathcal{G}$ as $B$ increases.
We conduct extensive experiments on three real-world datasets: MovieLens, DBLP, and Yelp. We aim to demonstrate the effectiveness of the optimizations in $\text{PHASE}_{\text{opt}}$ compared to PHASE, and to compare hypothesis-agnostic and hypothesis-aware samplers in terms of test significance, accuracy, and running time.
We first observe that $\text{PHASE}_{\text{opt}}$ is at least 20 times faster than PHASE. Then, by comparing $\text{PHASE}_{\text{opt}}$ with 11 state-of-the-art hypothesis-agnostic samplers, we find it achieves the best accuracy given the sampling budget across various hypothesis types and datasets. $\text{PHASE}_{\text{opt}}$ also exhibits fast convergence of hypothesis estimators and robust accuracy performance especially for hypotheses with longer paths or limited relevant nodes, edges, or paths in $\mathcal{G}$. For significance, $\text{PHASE}_{\text{opt}}$ consistently provides the most precise and reliable estimate, achieving the highest hypothesis significance among all samplers. Moreover, $\text{PHASE}_{\text{opt}}$ achieves high accuracy in the shortest amount of time, making it usable in practice.
2. Definitions
2.1. Attributed Graphs
Definition 0 (Attributed Graph).
An attributed graph is a directed graph $\mathcal{G}=(\mathcal{V,E})$, where $\mathcal{V}$ denotes the set of nodes and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges, represented by ordered pairs of nodes. It has a node type mapping function $\phi:\mathcal{V}\rightarrow\mathcal{T}$ and an edge type mapping function $\psi:\mathcal{E}\rightarrow\mathcal{R}$. Each node and edge has attributes.
Each node $v$ in $\mathcal{G}$ belongs to a specific node type $\phi(v)\in\mathcal{T}$, and each edge $e=(u,v)$ connecting nodes of type $\phi(u)$ to $\phi(v)$ belongs to a specific edge type $r=\psi(e)\in\mathcal{R}$. Once the relation $r$ exists, its inverse relation $r^{-1}$ naturally holds from $\phi(v)$ to $\phi(u)$. In this work, we assume each node in $\mathcal{G}$ has at least one incoming or outgoing edge. It implies the connectedness of the graph.
The DBLP Network in Figure 1a is an attributed graph containing four types of nodes and four types of edges. For example, paper nodes belong to the node type $\mathtt{paper}\in\mathcal{T}$ with five attributes, including title and year. The edge from paper to field of study (FOS) nodes is denoted by $\mathtt{WithDomain}\in\mathcal{R}$, with a weight indicating the paper’s relevance to a particular FOS. The reverse relationship $\mathtt{WithDomain}^{-1}\in\mathcal{R}$ also holds from FOS to paper nodes.
Definition 0 (Path).
A path $\mathcal{P}$ is defined as
$t_{1}\xrightarrow{r_{1}}\dots\xrightarrow{r_{l}}t_{l+1}$
where the node type $t_{i}$ and edge type $r_{i}$ can repeat; $l\geq 0$ is the length of $\mathcal{P}$. When $l=0$, $\mathcal{P}$ is a node and when $l=1$, it is an edge.
An attributed path is a path where each node has a list of attributes, referred to as a modifier. Figure 1b presents an example of a length-one attributed path, “conference papers written by Microsoft researchers”, and Figure 1c is an instance of that path.
Figure 2. The Sampling-based Hypothesis Testing Framework on attributed graphs.
2.2. Hypotheses on Attributed Graphs
We formally define path hypotheses on attributed graphs. Node and edge hypotheses are two special cases of path hypotheses.
Definition 0 (Path Hypothesis).
Given a path $\mathcal{P}=t_{1}\xrightarrow{r_{1}}t_{2}\xrightarrow{r_{2}}...\xrightarrow{r_{l}}t_{l+1}$ in $\mathcal{G}$, where $l\geq 1$, a path hypothesis is defined as:
[…]
In the DBLP network defined previously, co-authorship can be represented as a path:
$\mathcal{P}=\texttt{author}\xrightarrow{\mathtt{Authorship}^{-1}}\texttt{paper}\xrightarrow{\mathtt{Authorship}}\texttt{author}$
“The average citation of papers co-authored by Microsoft researchers is greater than 100” can be expressed as:
[…]
where $f_{\mathcal{P}}=\mathtt{paper}[\mathtt{citation}]$, ${\tt MSR}$ stands for microsoft research.
When $\mathcal{P}$ has length zero, a path hypothesis is reduced to a node hypothesis. An example of node hypothesis in the DBLP dataset can be “the average number of citations for conference papers is larger than 50” expressed as:
[…]
When $\mathcal{P}$ has length one, we refer it as an edge hypothesis. For instance, the edge hypothesis “the average $\mathtt{FOS\_weight}$ of conference papers on data mining is larger than 0.5” can be expressed as:
[…]
C (resp. DM) stands for conference papers (resp. DM data mining).
2.3. Problem Statement and Challenges
Given an attributed graph $\mathcal{G}$, our aim is to verify a node, edge, or path hypothesis $H$ and return true or false.
To address our problem, we formulate two objectives: O1 the hypothesis testing should be as accurate as possible; O2 the overall execution time should be as small as possible.
When testing a hypothesis such as “the average number of citations for conference papers is larger than 50”, a common approach includes collecting representative sample data and computing an aggregate value. In this case, the aggregate value corresponds to the average number of citations. However, directly collecting all relevant, i.e., satisfying the modifiers in the hypothesis, nodes, edges, or paths from $\mathcal{G}$ and computing an aggregate value can be time-consuming. Therefore, we adopt the commonly used approach of sampling the graph $\mathcal{G}$ into a smaller representative graph $\mathcal{S}$. This assumes a sampling budget $B$ that reflects the maximum size of $\mathcal{S}$. For the sake of simplicity, unless stated otherwise, we assume sampling an edge or a node requires the same unitary cost $1$.
A hypothesis estimator returns the aggregated value of nodes, edges, or paths in a hypothesis. The challenge of addressing O1 is that depending on the sampling method and the budget $B$, the hypothesis estimator obtained from $\mathcal{S}$ may differ from that of $\mathcal{G}$. We propose to first examine the applicability of existing sampling methods for graphs. Our observation is that the accuracy of those methods increases slowly as $B$ increases, resulting in longer times to achieve satisfactory accuracy. We conjecture that this is attributed to the lack of awareness of the underlying hypothesis during sampling. For instance, $\mathcal{S}$ should preserve as many relevant paths from $\mathcal{G}$ as possible for a path hypothesis. Hence, to address objective O2, we seek to design a sampling method that is aware of $H$ and preserves relevant nodes, edges, or paths in $H$. We further aim to demonstrate both theoretically and empirically that using our sampling, the hypothesis estimator on resulting $\mathcal{S}$ converges to its corresponding value in $\mathcal{G}$ as $B$ increases.
3. Sampling-based Hypothesis Testing
We propose a sampling-based hypothesis testing framework that supports both hypothesis-agnostic and hypothesis-aware graph samplers. After reviewing existing hypothesis-agnostic methods, we introduce our hypothesis-aware sampler, PHASE, in Section 3.2.1. It addresses objective O1 by ensuring convergence of hypothesis estimators, and objective O2 by incorporating hypothesis awareness into the sampling. Additionally, we implement two optimizations, $\text{PHASE}_{\text{opt}}$, to reduce the running time of PHASE in Section 3.2.2.
3.1. Hypothesis-Agnostic Samplers
Figure2 illustrates the sampling-based hypothesis testing framework we propose to test node, edge, and path hypotheses on attributed graphs. Based on whether the graph sampling method is hypothesis-agnostic or hypothesis-aware, there are two workflows indicated by dashed and dash-dot-dash arrows, respectively. Their distinction lies in where the hypothesis $H$ is introduced. In the hypothesis-aware workflow, $H$ is necessary during the sampling step, while in the hypothesis-agnostic workflow, $H$ is directly introduced before hypothesis testing.
Following dashed arrows from the left, the first procedure is to sample a graph $\mathcal{S}$ using a hypothesis-agnostic sampler, which requires $\mathcal{G}$ and $B$ only. The existing hypothesis-agnostic samplers can be categorised into three categories: node samplers that choose $B$ nodes from $\mathcal{G}$(Stumpf et al., 2005; Adamic et al., 2001), edge samplers that choose $B$ edges from $\mathcal{G}$(Krishnamurthy et al., 2005), and random walk based samplers that pick edges and nodes through the process of random walks(Gjoka et al., 2010; Lee et al., 2012; Leskovec and Faloutsos, 2006; Li et al., 2019). The second procedure is to conduct the hypothesis testing using $\mathcal{S}$ and $H$. Specifically, we extract all relevant nodes, edges, or paths from $\mathcal{S}$ and compute the required aggregated values for hypothesis testing. Finally, the acceptance or rejection result, p-value, and confidence interval are returned.
3.2. Hypothesis-Aware Samplers
3.2.1. PHASE Algorithm
We introduce PHASE, our Path-Hypothesis-Aware SamplEr (Algorithm1). PHASE aims to preserve nodes, edges, or paths specified in a hypothesis $H$. It can be integrated with any random walk based sampler that maintains the path-preserving property. To ease understanding, we illustrate it with FrontierS(Ribeiro and Towsley, 2010). FrontierS picks a node from $m$ random seeds based on degree-proportional probability and performs a random walk by uniformly selecting a neighbor to proceed. It repeats the process until reaching the sampling budget $B$.
Algorithm 1 PHASE
1:Input: $\mathcal{G}=(\mathcal{V,E})$, $B$, $H$
2:Output: a sampled graph $\mathcal{S}$
3:Initialize $L=(v_{1},v_{2},\ldots,v_{m})$ with $m$ randomly chosen nodes
4:$L_{w}=\text{AssignWeight}(L,H)$triangleright Assign $w_{h}$ to $x_{1}$ nodes and $w_{l}$ to other nodes.
5:$\mathcal{V}_{\mathcal{S}}=\{\}$, $\mathcal{E}_{\mathcal{S}}=\{\}$
6:while $B>m$ do
7:    Normalize $L_{w}$
8:    Select $v\in L$ with probability $L_{w}$
9:    $N\leftarrow N[v]$triangleright $N[v]$ is the set of neighbors of $v$.
10:    $N_{w}=\text{AssignWeight}(N,H)$triangleright Assign weights to nodes in $N$ according to the transition probability matrices.
11:    Select $u\in N$ with probability $N_{w}$
12:    $\mathcal{V}_{\mathcal{S}}$.update($v,u$)
13:    Replace $v$ by $u$ in $L$ and update $L_{w}$
14:    $B=B-1$
15:end while
16:return $\mathcal{S}=\{\mathcal{V}_{\mathcal{S}},\mathcal{E}_{\mathcal{S}}\}$
PHASE takes $\mathcal{G}$, $B$, and $H$ as inputs. For a node, edge, or path hypothesis, $H$ contains $\mathcal{P}$ with lengths zero, one, or more than one, respectively. It initializes $m$ seed nodes randomly in a list $L$ (Line 1), which increases the probability of selecting relevant nodes, edges, and paths from $\mathcal{G}$. Each seed is assigned a weight in $L_{w}$ (Line 2) to guide the selection of seed nodes for random walks later. These weights are determined based on heuristics, with nodes satisfying the first node modifier on $\mathcal{P}$ (denoted by $x_{1}$) assigned a higher weight $w_{h}$ and all other nodes assigned a lower weight $w_{l}$.
Lines 4-14 describe an iterative random walk to select nodes until $B$ is reached. During each iteration, the algorithm chooses a node $v$ from $L$ with probability $L_{w}$ (Line 6). Then, it selects one of $v$’s neighbors $u$ using a weighted selection process (Lines 7-9). Specifically, in Line 9, we adjust the transition probability to bias the random walks towards nodes, edges, or paths that satisfy $\mathcal{P}$ in $H$. The transition probability matrices for node, edge, and path hypotheses are depicted in Figures 3a, 3b, and 3c, respectively.
After adding nodes $u$ and $v$ to $\mathcal{V}_{S}$ (Line 10), the algorithm updates $L$ by replacing $v$ with $u$ (Lines 11) and decreases the sampling budget by one (Lines 12). This iteration continues until $B>m$. The resulting sampled graph $\mathcal{S}$ is the induced subgraph from $\mathcal{V}_{S}$.
$x_{1}$$y$$x_{1}$$w_{h}$$w_{l}$$y$$w_{h}$$w_{l}$(a)$x_{1}$$x_{2}$$y$$x_{1}$$w_{l}$$w_{h}$$w_{l}$$x_{2}$$w_{h}$$w_{l}$$w_{l}$$y$$w_{h}$$w_{l}$$w_{l}$(b)$x_{1}$$x_{2}$$x_{3}$$y$$x_{1},x_{1}$$w_{l}$$w_{h}$$w_{l}$$w_{l}$$x_{1},x_{2}$$w_{l}$$w_{l}$$w_{h}$$w_{l}$$x_{1},x_{3}$$w_{h}$$w_{l}$$w_{l}$$w_{l}$$x_{1},y$$w_{h}$$w_{l}$$w_{l}$$w_{l}$$\vdots$(c)
Figure 3. Transition probability matrices for node (a), edge (b), and path ($l=2$) (c) hypotheses, where $w_{h}\geq w_{l}>0$ denote transition probabilities. $x_{i}$ represents nodes in $\mathcal{G}$ satisfying the $i$-th node modifier on $\mathcal{P}$, while $y$ represents nodes not satisfying any node modifier on $\mathcal{P}$. (a) and (b) involve 1st-order random walks, whereas (c) involves 2nd-order random walks because the probability of selecting the next node depends on both the current and previous nodes.
3.2.2. $\text{PHASE}_{\text{opt}}$ Algorithm
We further propose optimizations to reduce execution time. Algorithm 2 introduces two lines that replace line 7 in Algorithm 1. Firstly, in densely connected graphs, the computation of neighbor weights in step 8 of Algorithm 1 can be time-consuming due to the potentially large number of neighbors. To mitigate this, we randomly sample a subset of $\min\{|N|,n\}$ neighbors as candidates for selection, where $n$ is a parameter. Secondly, we adopt a non-backtracking approach by avoiding the selection of visited nodes during the sampling process(Lee et al., 2012). This strategy prevents cycles and unnecessary node visits.
Algorithm 2 $\text{PHASE}_{\text{opt}}$
(showing only 2 lines that replace line 7 in Algorithm 1
1:$N^{\prime}\leftarrow N[v]-\mathcal{V_{S}}$ (Optim 2)
2:$N\leftarrow$ Select $\text{min}\{|N^{\prime}|,n\}\text{ nodes randomly from }N^{\prime}$ (Optim 1)
Complexity. On average, each node has $2|\mathcal{E}|/|\mathcal{V}|$ neighbors to be examined by PHASE, resulting in a time complexity of $O(B\times 2|\mathcal{E}|/|\mathcal{V}|)$. On the other hand, by constraining the number of neighbors, $\text{PHASE}_{\text{opt}}$ achieves a time complexity of $O(B)$.
3.2.3. Convergence of Hypothesis Estimators
Let us demonstrate the convergence of hypothesis estimators using our proposed sampling algorithms (Challenge 1).
By Lemma 5.1 from(Ribeiro and Towsley, 2010), FrontierS is equivalent to the sampling process of a SRW over $\mathcal{G}^{m}=(\mathcal{V}^{m},\mathcal{E}_{m})$, where $\mathcal{G}^{m}$ is the $m$-th Cartesian power of $\mathcal{G}$ and
$\mathcal{V}^{m}=\{(v_{1},\ldots,v_{m})|v_{1}\in\mathcal{V}\wedge\dots\wedge v_{m}\in\mathcal{V}\}$
is the $m$-th Cartesian power of $\mathcal{V}$. $\forall\boldsymbol{u,v}\in\mathcal{V}^{m}$, $(\boldsymbol{u,v})\in\mathcal{E}_{m}$ if there exists $i$ such that $(u_{i},v_{i})\in\mathcal{E}$ and for all $j\neq i$, $u_{j}=v_{j}$. Given $\mathcal{G}$ is a directed, connected, and non-bipartite graph, in the steady state, the sampled edges of FrontierS form a stationary sequence and satisfy the following Strong Law of Large Numbers (SLLN), the proof of which can be found in (Ribeiro and Towsley, 2010).
Theorem 1 (SLLN).
For any function $f$, where
$\sum_{(u,v)\in\mathcal{E}}|f(u,v)|<\infty$,
$\lim_{B\to\infty}\frac{1}{B}\sum^{B}_{i=1}f(u_{i},v_{i})\to\frac{1}{|E|}\sum_{\forall(u,v)\in\mathcal{E}}f(u,v)$
almost surely, i.e. the event occurs with probability one.
For a path hypothesis with length $l$, it involves $(l-1)$-th order random walk, which is equivalent to a SRW over $\mathcal{G}^{l-1}=(\mathcal{V}^{l-1},\mathcal{E}_{l-1})$(Lovász, 1993). Hence, PHASE’s random walks can be viewed as a SRW over $\mathcal{G}^{l-1+m}=(\mathcal{V}^{l-1+m},\mathcal{E}_{l-1+m})$. As PHASE ensures that the resulting random walks over $\mathcal{G}^{l-1+m}$ maintain both irreducibility and aperiodicity, it preserve stationarity and adhere to the SLLN. $\text{PHASE}_{\text{opt}}$ also follows the SLLN because the stationary distribution of the NBRW remains the same as that of the SRW.
By Theorem1, we can construct hypothesis estimators for node, edge, and path hypotheses and prove their convergence. Here, we focus on scenarios where agg represents an average function. The hypothesis estimators for other aggregation functions, such as maximum and minimum, can be derived in a similar manner.
A path hypothesis ($l\geq 2$): $\text{avg}(f_{\mathcal{P}}|M_{t_{i}},\forall t_{i}\text{ on }\mathcal{P})$ has a primary subject $f_{\mathcal{P}}$. Let $\mathcal{P}^{*}$ be all relevant paths in $\mathcal{G}$, and $\mathcal{P_{G}}$ be all paths with the same length as $\mathcal{P}$ in $\mathcal{G}$. The mean value of the path hypothesis, $\theta_{path}$, is
(1) $\theta_{path}=\frac{1}{|\mathcal{P}^{*}|}\sum_{\forall\mathcal{P}\in\mathcal{P_{G}}}T(\mathcal{P})$
where $T(\mathcal{P})=f_{\mathcal{P}}\times\mathbbm{1}_{M_{t_{i}}\subseteq\mathcal{L}_{\phi(t_{i})}\forall t_{i}\text{ on }\mathcal{P}}$. Replacing $\mathcal{G}$ with the sampled graph $\mathcal{S}$, the estimator for $\theta_{path}$ is
(2) $\hat{\theta}_{path}=\frac{1}{|\mathcal{P}^{*}\cap\mathcal{P_{S}}|}\sum_{\forall\mathcal{P}\in\mathcal{P_{S}}}T(\mathcal{P})$
When $B$ goes to infinity, $\mathcal{S}$ converges to $\mathcal{G}$. By Theorem1, we have
(3) $\lim_{B\to\infty}\frac{1}{|\mathcal{P}^{*}\cap\mathcal{P_{S}}|}\sum_{\forall\mathcal{P}\in\mathcal{P_{S}}}T(\mathcal{P})\to\frac{1}{|\mathcal{P}^{*}|}\sum_{\forall\mathcal{P}\in\mathcal{P_{G}}}T(\mathcal{P})$
almost surely. Hence, $\hat{\theta}_{path}$ is an asymptotically unbiased estimator of $\theta_{path}$. Hypothesis estimators for node and edge hypotheses can be derived similarly.
4. Experiments
The goal of our experiments is twofold: 1) test the effectiveness of optimizations in $\text{PHASE}_{\text{opt}}$ compared to PHASE (Section 4.3), and 2) compare hypothesis-agnostic and hypothesis-aware samplers in terms of test significance (Section 4.4), accuracy (Section 4.5), and time (Section 4.6). The source code is publicly available  [^1] https://github.com/Carrieww/GraphHT.
4.1. Experimental Setup
Table 1. Statistics of datasets
Dataset #(Nodes) #(Edges) Density #(Node Types) #(Edge Types)
MovieLens 9,705 996,656 1.06e-02
2
1
DBLP 1,623,013 11,040,170 4.19e-06
4
4
Yelp 2,136,118 6,743,879 1.48e-06
2
1
Datasets. We used three datasets (Harper and Konstan, 2016; Tang et al., 2008), extracted from real attributed networks: MovieLens, DBLP, and Yelp. Table 1 shows the statistics of the datasets.
Samplers We compare $\text{PHASE}_{\text{opt}}$ with eleven existing samplers:
  • Node samplers: Random Node Sampler (RNS), Degree-Based Sampler (DBS)
  • Edge samplers: Random Edge Sampler (RES)
  • Random walk based samplers: Simple Random Walk (SRW), Frontier Sampler (FrontierS), Non-Backtracking Random Walk (NBRW), Random Walk with Restarter (RWR), Metropolis-Hastings Random Walk (MHRW), Snaw Ball Sampler (SBS), Forest Fire Sampler (FFS), Shortest Path Sampler (ShortestPathS)
Hypotheses. For each dataset (DB) and hypothesis type (N for node hypotheses), we provide three example hypotheses (labeled as DB-N1) in experiments. For example, on Yelp, we explore an edge hypothesis testing if “fast food average ratings exceed 4”. In DBLP, we examine a path hypothesis to see if “papers co-authored by Chinese institutes on data mining have over 50 citations”. To assess path length effects, we further test two path hypotheses with lengths of three and four in DBLP.
Parameter Choice. For PHASE and $\text{PHASE}_{\text{opt}}$, We set $m=50$ to maintain the path-preserving capability of multi-dimensional random walks, and $n=30$ to strike a balance between accuracy and time efficiency. Additionally, we choose $w_{h}=10$ and $w_{l}=0.1$ to regulate the bias towards relevant nodes, edges, and paths. The sampling budget $B$ is maintained as a proportion of the total number of nodes in $\mathcal{G}$ for all samplers. As for existing hypothesis-agnostic samplers, we use the best parameters in their respective settings. We report an average of 30 runs for every evaluation measure.
4.2. Evaluation Measures
Accuracy. Accuracy at a specific sampling proportion reflects the effectiveness of a sampling method. The accuracy below measures the number of matched hypothesis testing results on $\mathcal{G}$ and $\mathcal{S}$:
$\text{Accuracy}=\frac{1}{k}\sum_{k}\mathbbm{1}_{H(\mathcal{G})==H(\mathcal{S})}$
where $H(\mathcal{G})$ and $H(\mathcal{S})$ return 0 (false) or 1 (true).
Time. We measure the total execution time, including both the sampling time and the time taken to extract relevant information from $\mathcal{S}$ for hypothesis testing.
Table 2. Average execution time (sec) of PHASE and $\text{PHASE}_{\text{opt}}$ for DBLP node, edge, and path hypotheses.
Node Edge Path
PHASE 115.66 539.15 441.3
$\text{PHASE}_{\text{opt}}$ 5.56 8.76 5.91
Table 3. The accuracy of eleven existing samplers and $\text{PHASE}_{\text{opt}}$ on three datasets and three types of hypothesis.
Dataset
Hypothesis Type
Sampling Proportion (%)
$\text{PHASE}_{\text{opt}}$ RES RNS DBS SRW NBRW RWR MHRW ShortestPathS FrontierS FFS SBS
Node
1
0.9 0.89 0.83 0.86 0.88 0.89 0.87 0.89 0.86 0.84 0.83 0.56
Edge
2.5
0.98 0.68 0.77 0.98 0.99 1 0.99 0.98 0.99 0.73 0.93 0.91
MovieLens
Path
5
0.99 0.1 0.88 0.83 0.82 0.89 0.91 0.98 0.95 0.38 0.85 0.62
Node
0.2
0.96 0.48 0.87 0.93 0.92 0.91 0.9 0.94 0.92 0.92 0.94 0.88
Edge
0.2
0.76 0.48 0 0.71 0.73 0.69 0.7 0.29 0.69 0.63 0.7 0.42
DBLP
Path
0.2
0.89 0 0 0 0.26 0.3 0.32 0.18 0.33 0.043 0.3 0.12
Node
0.1
0.99 0.65 0.77 0.61 0.69 0.69 0.69 0.77 0.7 0.64 0.66 0.48
Edge
1
1 0.73 0.54 0.76 0.79 0.91 0.87 0.84 0.76 0.77 0.79 0.71
Yelp
Path
1
0.99 0.11 0.05 0.79 0.78 0.78 0.72 0.8 0.99 0.42 0.67 0.42
Table 4. The execution time (sec) of eleven existing samplers and $\text{PHASE}_{\text{opt}}$ on three datasets and three types of hypothesis.
Dataset
Hypothesis Type
Sampling Proportion (%)
$\text{PHASE}_{\text{opt}}$ RES RNS DBS SRW NBRW RWR MHRWS ShortestPathS FrontierS FFS SBS
MovieLens
Node
1
0.083 0.99 0.023 0.077 0.057 0.06 0.05 0.06 0.063 0.083 0.067 0.047
Edge
2.5
0.45 0.99 0.11 0.36 0.31 0.36 0.33 0.29 0.26 0.33 0.35 0.31
Path
5
4.92 1.03 0.34 4.32 4.53 4.55 4.76 3.10 0.95 0.38 3.87 3.08
DBLP
Node
0.2
5.56 18.70 0.55 6.98 7.98 8.22 9.67 1.66 31.32 10.48 5.73 3.03
Edge
0.2
8.76 22.57 0.90 10.07 14.76 12.89 12.32 3.53 33.91 14.46 8.87 5.27
Path
0.2
5.44 19.61 0.71 6.87 8.24 8.39 8.13 2.17 31.13 9.37 5.33 3.01
Yelp
Node
0.1
2.56 13.84 1.02 6.81 1.19 0.96 1.16 1.12 1.84 1.68 1.02 6.58
Edge
1
19.42 16.88 2.01 13.61 8.58 8.97 9.17 6.93 30.23 10.22 10.27 6.03
Path
1
56.97 14.35 1.53 23.48 15.97 16.61 19.03 7.35 37.81 9.95 25.61 19.94
4.3. PHASE vs $\text{PHASE}_{\text{opt}}$
We evaluate the effectiveness of optimizations in $\text{PHASE}_{\text{opt}}$ compared to PHASE using the DBLP dataset, as shown in Table 2. $\text{PHASE}_{\text{opt}}$ is at least 20 times faster than PHASE while maintaining comparable accuracy (less than 5% accuracy reduction). Consequently, we exclusively use $\text{PHASE}_{\text{opt}}$ for subsequent experiments.
4.4. Significance
(a) P-value Plot
(b) CI Plot
Figure 4. DBLP p-value and CI plot for the hypothesis DB-P1
To examine statistical significance and estimation precision, we analyze the trend of p-values and confidence intervals (CI) with increasing $B$. We illustrate this trend using a path hypothesis on DBLP as all other hypotheses show similar trends. As shown in Figure 3(a), the p-value of $\text{PHASE}_{\text{opt}}$ consistently stays below the significance level (e.g., 0.05) and decreases gradually. This trend indicates stronger evidence against the null hypothesis, showing higher statistical significance as $B$ increases.
Figure 3(b) shows narrowing CI with increasing $B$, indicating more precise and reliable estimations of the hypothesis estimator. $\text{PHASE}_{\text{opt}}$ consistently achieves the narrowest bound at any $B$, indicating the most precise and reliable estimate among all samplers.
(a) Accuracy ($l=3$)
(b) Accuracy ($l=4$)
Figure 5. Accuracy versus sampling proportion plots for DBLP path hypotheses with length three (a) and four (b).
4.5. Accuracy
Table 3 presents the accuracy performance of eleven existing samplers and $\text{PHASE}_{\text{opt}}$ at a specific sampling proportion, shown in column three. We select this proportion to balance stability and discriminative power, ensuring the accuracy is sufficiently stabilized without saturating across all samplers. For each dataset and type of hypothesis, the accuracy is the average accuracy of three hypothesis examples, each derived from an average of 30 runs. The highest accuracy in each row is highlighted in bold.
In Table 3, $\text{PHASE}_{\text{opt}}$ demonstrates robust performance in all cases, except for MovieLens edge hypotheses, in where $\text{PHASE}_{\text{opt}}$ slightly trails behind NBRW. Comparing among columns, some random walk based samplers, namely SRW, RWR, and ShortestPathS, exhibit good accuracy performance for edge and path hypothesis. On the contrary, RES, RNS, and DBS show ineffectiveness in handling path hypotheses. These observations align with their sampling mechanisms. These node and edge samplers can hardly preserve the path information from $\mathcal{G}$ in the sampled graph. Additionally, comparing $\text{PHASE}_{\text{opt}}$ and FrontierS, $\text{PHASE}_{\text{opt}}$’s superior performance indicates the efficacy of the two weight functions in enhancing the sampler’s hypothesis-awareness to sample relevant nodes, edges, and paths.
(a) Time $(l=3)$
(b) Time ($l=4$)
Figure 6. Time versus accuracy plots for DBLP path hypotheses with length three (a) and four (b).
(a) DB-N1
(b) DB-N2
(c) DB-N3
(d) DB-N3
(e) DB-E1
(f) DB-E2
(g) DB-E3
(h) DB-E3
(i) DB-P1
(j) DB-P2
(k) DB-P3
(l) DB-P3
Figure 7. Comparison of the top five sampling methods for accuracy and time (d, h, l) efficiency across three node (a, b, c), three edge (e, f, g), and three path hypotheses (i, j, k) for the DBLP dataset.
We establish an accuracy ranking for all samplers by averaging their accuracy performance per column in Table 3 and select the top five: $\text{PHASE}_{\text{opt}}$, NBRW, ShortestPathS, RWR, and SRW. These samplers’ accuracy convergence on individual hypotheses is compared in Figures 7 and 8 for DBLP and Yelp datasets, respectively. The figure for MovieLens is excluded due to space constraints and the similarity of results. In these figures, a-c, e-g, and i-k display the accuracy performance of three node, three edge, and three path hypotheses, respectively. Hypotheses of the same type are structured to have progressively fewer relevant nodes, edges, or paths in $\mathcal{G}$. The x-axis is truncated to clearly show the convergence.
In Figures 7 and 8, $\text{PHASE}_{\text{opt}}$ consistently outperforms other methods across all sampling proportions for most hypotheses. For hypotheses of the same type, $\text{PHASE}_{\text{opt}}$, with hypothesis-awareness, consistently demonstrates robust accuracy performance for the hypotheses with fewer relevant nodes, edges, or paths in $\mathcal{G}$. This is particularly evident for DBLP edge and path hypotheses. When relevant nodes, edges, and paths in $\mathcal{G}$ are abundant, conventional sampling methods may have sampled a sufficient number of relevant nodes, edges, or paths to achieve high accuracy, which can diminish the distinctive advantage of $\text{PHASE}_{\text{opt}}$.
Delving into Figure 7, $\text{PHASE}_{\text{opt}}$ dominates in node and path hypotheses for the DBLP dataset. However, ShortestPathS slightly outperforms $\text{PHASE}_{\text{opt}}$ in edge hypotheses one and two, possibly due to the high betweenness centrality of these edges, implying that many shortest paths traverse those edges. Moving to Figure 8, $\text{PHASE}_{\text{opt}}$ demonstrates clear superiority across all types of hypotheses for the Yelp dataset. Nonetheless, ShortestPathS occasionally exhibits competitive performance compared to $\text{PHASE}_{\text{opt}}$ in path hypotheses, which may also be attributed to the betweenness centrality of those paths in $\mathcal{G}$.
Figure 5 shows how path length affects accuracy using the DBLP dataset. $\text{PHASE}_{\text{opt}}$ enables faster convergence of accuracy with longer path lengths, outperforming other samplers.
4.6. Scalability
(a) YP-N1
(b) YP-N2
(c) YP-N3
(d) YP-N3
(e) YP-E1
(f) YP-E2
(g) YP-E3
(h) YP-E3
(i) YP-P1
(j) YP-P2
(k) YP-P3
(l) YP-P3
Figure 8. Comparison of the top five sampling methods for accuracy and time (d, h, l) efficiency across three node (a, b, c), three edge (e, f, g), and three path hypotheses (i, j, k) for the Yelp dataset.
In Table 4, samplers’ execution time increases as the dataset becomes larger. Among them, RNS generally exhibits the shortest execution time as it uniformly samples nodes randomly. The performance of $\text{PHASE}_{\text{opt}}$ varies across datasets. For the Yelp dataset, it ranks among the highest, particularly for the path hypothesis. This can be attributed to the limited node types in Yelp, resulting in longer extraction times for condensed relevant paths from $\mathcal{S}$. For the DBLP dataset, $\text{PHASE}_{\text{opt}}$ ranks among the lowest five. However, instead of considering execution time alone, we should consider both execution time and accuracy performance to evaluate a sampler’s efficiency and effectiveness.
For the top five samplers ranked by their accuracy performance, we plot their accuracy versus execution time in subfigures d, h, and l of Figures 7 and 8. Due to space limitations, only the third hypothesis, which consists of the least number of relevant nodes, edges, or paths in the input graph, of each type is included. First of all, we observe that the curve of $\text{PHASE}_{\text{opt}}$ is always at the bottom among others. This means that $\text{PHASE}_{\text{opt}}$ consistently achieves high accuracy in the shortest amount of time across all plots. Additionally, its execution time does not exhibit exponential growth compared to other existing hypothesis-agnostic samplers. This can be attributed to $\text{PHASE}_{\text{opt}}$’s hypothesis-awareness, which allows it to efficiently sample relevant nodes, edges, and paths from the input graph to verify the hypothesis.
The impact of path length on accuracy and execution time is illustrated in Figure 6. As the path length increases, $\text{PHASE}_{\text{opt}}$ consistently achieves the highest accuracy within the shortest execution time.
5. Related Work
There has been a lot of research interest on hypothesis testing (Hunter et al., 2008; Gibbs et al., 2022; Arias-Castro and Verzelen, 2014; Xia and Li, 2017; Croft et al., 2011; Ghoshdastidar and von Luxburg, 2018; Omidvar-Tehrani et al., 2019), proposing sampling techniques on graphs (Leskovec and Faloutsos, 2006; Lee et al., 2012; Li et al., 2019). In this section we summarize the most representative works.
Hypothesis Testing on Graphs. Hypotheses on graphs can be categorized by their object of interest (Zhao et al., 2011; Arias-Castro and Verzelen, 2014; Bickel and Sarkar, 2013). Tang et al. and Ghoshdastidar et al. (Tang et al., 2017; Ghoshdastidar et al., 2017, 2020) extend the one-sample problem into two-sample to test whether two groups of random graphs are obtained by the same generative model or with the same graph distribution. In (Xia and Li, 2017), nodes are objects of interest, and one-sample hypothesis testing is used to detect conditional dependence between nodes in brain connectivity networks.
Our goal is more related to works when the object of interest is nodes. However, there are two differences between our work and the existing ones. First, we enable more expressive hypotheses, including edge and path hypothesis, on graphs. Second, we leverage graph sampling methods on graph hypothesis testing.
Sampling Methods on Graphs. Most sampling methods are designed to sample a representative subgraph to accurately estimate graph properties (Leskovec and Faloutsos, 2006; Stutzbach et al., 2009; Maiya and Berger-Wolf, 2010; Hübler et al., 2008). Leskovec and Faloutsos (Leskovec and Faloutsos, 2006) are the first to study this problem in real-world networks (Leskovec and Faloutsos, 2006) by proposing sampling techniques to maintain degree distribution, clustering coefficient and distribution of component sizes. Hübler et al. (Hübler et al., 2008) propose the Metropolis-Hastings sampling method. Later, Maiya and Berger-Wolf (Maiya and Berger-Wolf, 2010) propose a community representativeness sample and design a community structure expansion sampler. Many sampling methods are proposed to improve the efficiency and convergence rate of simple random walk (Lee et al., 2012; Li et al., 2019). Besides representative sampling, other sampling methods are designed for specific tasks, such as graph compression (Besta et al., 2019), community detection(Maiya and Berger-Wolf, 2010), and graph visualization (Rafiei and Curial, 2005; Wu et al., 2017). However, none of them is designed for hypothesis testing on attributed graphs.
6. Conclusion
In this paper, we developed a framework for hypothesis testing on large attributed graphs. Our framework accommodates eleven common hypothesis-agnostic samplers as well as new hypothesis-aware samplers. We proposed dedicated optimizations to speed up sampling while achieving high test significance and accuracy. We demonstrated theoretically and empirically that the convergence rate of our hypothesis-aware sampling.
Our work opens several new directions in the area of hypothesis sampling on graphs. The first direction is to examine additional optimizations that make use of domain-specific information on the input graph. The second direction is to handle more expressive hypotheses and specify two-sample and multiple sample scenarios.

### v2

@@
-A Sampling-based Framework for Hypothesis Testing on Large Attributed Graphs [Scalable Data Science]
+A Sampling-based Framework for Hypothesis Testing on Large Attributed Graphs
 Abstract.
 Hypothesis testing is a statistical method used to draw conclusions about populations from sample data, typically represented in tables.
-With the prevalence of graph representations in real-life applications, hypothesis testing in graphs is gaining importance.
-In this work, we formalize node, edge, and path hypotheses in attributed graphs.
+With the prevalence of graph representations in real-life applications, hypothesis testing on graphs is gaining importance.
+In this work, we formalize node, edge, and path hypotheses on attributed graphs.
 We develop a sampling-based hypothesis testing framework, which can accommodate existing hypothesis-agnostic graph sampling methods.
-To achieve accurate and efficient sampling, we then propose a Path-Hypothesis-Aware SamplEr, PHASE, an $m$-dimensional random walk that accounts for the paths specified in a hypothesis.
+To achieve accurate and time-efficient sampling, we then propose a Path-Hypothesis-Aware SamplEr, PHASE, an $m$-dimensional random walk that accounts for the paths specified in the hypothesis.
 We further optimize its time efficiency and propose $\text{PHASE}_{\text{opt}}$.
-Experiments on real datasets demonstrate the ability of our framework to leverage common graph sampling methods for hypothesis testing, and the superiority of hypothesis-aware sampling in terms of accuracy and time efficiency.
+Experiments on three real datasets demonstrate the ability of our framework to leverage common graph sampling methods for hypothesis testing, and the superiority of hypothesis-aware sampling methods in terms of accuracy and time efficiency.
+PVLDB Reference Format:
+PVLDB, 17(11): 3192 - 3200, 2024.
+doi:10.14778/3681954.3681993
+This work is licensed under the Creative Commons BY-NC-ND 4.0 International License.
+Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license.
+For any use beyond those covered by this license, obtain permission by emailing info@vldb.org.
+Copyright is held by the owner/author(s).
+Publication rights licensed to the VLDB Endowment.
+Proceedings of the VLDB Endowment, Vol. 17, No. 11 ISSN 2150-8097.
+doi:10.14778/3681954.3681993
+PVLDB Artifact Availability:
+The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣sethttps://github.com/Carrieww/GraphHT.
 1. Introduction
 Hypothesis testing finds widespread application in various domains such as marketing, healthcare, and social science (Newey and McFadden, 1994).
-Hypotheses are usually tested on representative samples since it is impractical or even impossible to extract data from entire populatations due to size, accessibility or cost.
+Hypotheses are usually tested on representative samples since it is impractical, or even impossible to extract data from entire populations due to size, accessibility or cost.
 For instance, snowball sampling has been proven effective in accessing a hidden population like drug abusers through a chain-referral mechanism  (Waters, 2015; Goodman, 1961).
 In this paper, we study hypothesis testing on graphs.
-Graphs can be used to represent many real-world applications such as social, bibliographic and transportation networks where entities are nodes and edges reflect relationships between entities.
-Specifically, we work on Attributed Graphs that are graphs with multiple types of nodes and edges, both of which are labeled with attributes containing valuable information for graph analytics (Croft et al., 2011; Omidvar-Tehrani et al., 2019).
-In a bibliographic network like DBLP (see Figure 1a and Figure 1b for an example and an attributed path of length one, respectively), hypotheses expressed on nodes, edges, and paths unveil interesting aspects of collaboration behavior and research trends based on citation evolution over time.
-Specifically, a node hypothesis can be “the average citation of conference papers is greater than average”, and a path hypothesis can be “the minimum number of citations of conference papers written by Microsoft researchers is greater than average”.
-Similarly, on social platforms like Yelp, hypothesis testing can be used to check how different users review a product and study its adoption.
 Figure 1. DBLP network schema and paths
+Graphs can represent real-world applications such as social, bibliographic, and transportation networks where entities are nodes and edges reflect relationships between them.
+We focus on attributed graphs that contain multiple types of nodes and edges, each labeled with attributes that provide valuable information for graph analytics (Croft et al., 2011; Omidvar-Tehrani et al., 2019).
+In a bibliographic network like DBLP (see Figure 1a), hypotheses expressed on nodes, edges, and paths unveil interesting aspects of collaboration behavior and research trends based on citation patterns.
+A node hypothesis examines attributes of a single node type.
+For example, “the average citation of conference papers is greater than average” concerns the citation attribute of the conference paper nodes.
+Path hypotheses consider attributes along a defined path, such as a path connecting authors from Microsoft to their conference papers (see Figure 1b and 1c for the path and an path instance, respectively).
+An example of path hypothesis is “the minimum number of citations of conference papers written by Microsoft researchers is greater than average”.
 Objectives and Challenges.
-Given an input attributed graph $\mathcal{G}$, our aim is to verify a hypothesis $H$ and return true or false.
-We hence seek to achieve two objectives: O1 - hypothesis testing should be as accurate as possible; O2 - execution time should be as small as possible.
+Given an attributed graph $\mathcal{G}$, our aim is to verify a hypothesis $H$ and return a true or false result.
+We aim to achieve two objectives: O1 - ensure hypothesis testing is as accurate as possible; O2 - minimize execution time.
 To the best of our knowledge, there is no existing literature addressing hypotheses on attributed graphs.
-A straightforward approach is to obtain a sampled graph $\mathcal{S}$ from $\mathcal{G}$ using hypothesis-agnostic graph sampling methods (Maiya and Berger-Wolf, 2010; Li et al., 2019; Adamic et al., 2001).
-However, the challenge is that depending on the sampling method and on the budget $B$, i.e., the size of $\mathcal{S}$, the hypothesis estimator, which computes aggregated values of nodes, edges, or paths specified in the hypothesis may return different values on $\mathcal{S}$ and $\mathcal{G}$.
+A straightforward approach is to sample a subgraph $\mathcal{S}$ from $\mathcal{G}$ using hypothesis-agnostic sampling methods (Maiya and Berger-Wolf, 2010; Li et al., 2019; Adamic et al., 2001).
+Sampling is important when the full graph is not accessible or when the hypothesis targets very specific nodes, edges, or paths that are time-consuming to collect.
+However, hypothesis-agnostic samplers struggle to achieve both objectives.
+First, hypothesis-agnostic samplers may miss relevant nodes, edges, or paths, i.e., those requested by the hypothesis, especially when the sampling budget $B$, i.e., the size of $\mathcal{S}$, is small.
+The irrelevant nodes, edges, or paths, which must be filtered out for hypothesis testing, can compromise accuracy.
 This raises the accuracy challenge behind O1.
-Additionally, as $B$ increases, the accuracy of the resulting hypotheses increases.
-However, the speed at which the accuracy converges highly depends on the sampling.
-This raises the efficiency challenge behind O2.
+We show in Section 3.2.3 that the hypothesis estimator, which computes aggregate values of nodes, edges, or paths specified in $H$, converges to the ground truth in $\mathcal{G}$ as $B$ increases, resulting in accuracy approaching one.
+However, the rate of convergence depends on the sampler, presenting a time efficiency challenge behind O2.
+By optimizing the sampler design, we aim to balance accuracy and time efficiency, enabling earlier sampling halts with accurate hypothesis testing results.
 Contributions.
-We classify hypotheses on attributed graphs into three types: node, edge, and path hypotheses.
-We develop a sampling-based hypothesis testing framework, which accommodates common hypothesis-agnostic sampling methods, such as random node sampler (Stumpf et al., 2005), simple random walk (Gjoka et al., 2010), and non-backtracking random walk (Lee et al., 2012).
-The lack of awareness of the input hypothesis in existing sampling methods may slow down the convergence speed of accuracy.
-Indeed, the sampled graph $\mathcal{S}$ should preserve as many relevant paths according to $H$ from $\mathcal{G}$ as possible for a path hypothesis.
+We classify hypotheses on attributed graphs into three types: node, edge, and path ones.
+We develop a sampling-based hypothesis testing framework, which accommodates common hypothesis-agnostic samplers, such as random node sampler (Stumpf et al., 2005), random walk (Gjoka et al., 2010), and non-backtracking random walk (Lee et al., 2012).
+The lack of awareness of the input hypothesis in existing sampling methods may slow down the rate of convergence of accuracy.
 Therefore, to address objectives O1 and O2, we design PHASE, a Path-Hypothesis-Aware SamplEr.
-PHASE is aware of $H$ and preserves its nodes, edges, or paths.
-Consequently, the resulting $\mathcal{S}$ is more likely to provide relevant information for testing the hypothesis.
+PHASE is aware of $H$ and preserves the corresponding nodes, edges, or paths.
+Consequently, the resulting $\mathcal{S}$ is more likely to accurately test the hypothesis.
 PHASE employs $m\geq 1$ dependent random walks with two weight functions to ensure path-hypothesis-awareness.
-One function biases the seed selection toward the first node in the path hypothesis.
-The other biases the transition probability towards the nodes specified in the hypothesis.
+One function prioritizes the seed selection toward the first node in the path hypothesis.
+The other steers the transition probability towards the nodes specified in the hypothesis.
 We further propose $\text{PHASE}_{\text{opt}}$ to improve the time efficiency of PHASE.
-$\text{PHASE}_{\text{opt}}$ reduces computation overhead through random sampling of neighbors.
-It also adopt a non-backtracking approach to prioritize the selection of relevant nodes, edges, and paths.
-We demonstrate both theoretically and empirically that the hypothesis estimator on $\mathcal{S}$ converges to its corresponding value in $\mathcal{G}$ as $B$ increases.
+$\text{PHASE}_{\text{opt}}$ adopts a non-backtracking approach to avoid selecting previously visited nodes.
+It also reduces computation overhead by fixing the number of neighbors to examine.
+We show both theoretically and empirically that, for all samplers, the hypothesis estimator converges to the ground truth as $B$ increases, and our proposed samplers ensures earlier and smoother convergence of the hypothesis estimator.
 We conduct extensive experiments on three real-world datasets: MovieLens, DBLP, and Yelp.
-We aim to demonstrate the effectiveness of the optimizations in $\text{PHASE}_{\text{opt}}$ compared to PHASE, and to compare hypothesis-agnostic and hypothesis-aware samplers in terms of test significance, accuracy, and running time.
-We first observe that $\text{PHASE}_{\text{opt}}$ is at least 20 times faster than PHASE.
-Then, by comparing $\text{PHASE}_{\text{opt}}$ with 11 state-of-the-art hypothesis-agnostic samplers, we find it achieves the best accuracy given the sampling budget across various hypothesis types and datasets.
-$\text{PHASE}_{\text{opt}}$ also exhibits fast convergence of hypothesis estimators and robust accuracy performance especially for hypotheses with longer paths or limited relevant nodes, edges, or paths in $\mathcal{G}$.
-For significance, $\text{PHASE}_{\text{opt}}$ consistently provides the most precise and reliable estimate, achieving the highest hypothesis significance among all samplers.
-Moreover, $\text{PHASE}_{\text{opt}}$ achieves high accuracy in the shortest amount of time, making it usable in practice.
+We aim to demonstrate the effectiveness of two optimizations in $\text{PHASE}_{\text{opt}}$ compared to PHASE, and to compare the test significance, accuracy, and execution time of hypothesis-agnostic and hypothesis-aware samplers.
+We observe $\text{PHASE}_{\text{opt}}$ is at least 43 times faster than PHASE with less than 4% accuracy difference in DBLP.
+For significance, $\text{PHASE}_{\text{opt}}$ consistently delivers the most precise and reliable estimates.
+Compared to 11 state-of-the-art hypothesis-agnostic samplers, we find $\text{PHASE}_{\text{opt}}$ excels in accuracy when $B$ is fixed across various hypothesis types and datasets.
+It also demonstrates robust accuracy performance, especially for difficult hypotheses, i.e., those with longer paths or fewer relevant nodes, edges, or paths in $\mathcal{G}$.
+Moreover, the high accuracy achieved in the shortest amount of time makes our proposed sampler usable in practice.
 2. Definitions
@@
 For example, paper nodes belong to the node type $\mathtt{paper}\in\mathcal{T}$ with five attributes, including title and year.
-The edge from paper to field of study (FOS) nodes is denoted by $\mathtt{WithDomain}\in\mathcal{R}$, with a weight indicating the paper’s relevance to a particular FOS.
-The reverse relationship $\mathtt{WithDomain}^{-1}\in\mathcal{R}$ also holds from FOS to paper nodes.
+Edges from paper to field of study (FOS) nodes belong to the edge type $\mathtt{WithDomain}\in\mathcal{R}$ with a weight attribute indicating the paper’s relevance to an FOS.
+The reverse relationship $\mathtt{WithDomain}^{-1}\in\mathcal{R}$ holds accordingly from FOS to paper nodes.
 Definition 0 (Path).
@@
 where the node type $t_{i}$ and edge type $r_{i}$ can repeat; $l\geq 0$ is the length of $\mathcal{P}$. When $l=0$, $\mathcal{P}$ is a node and when $l=1$, it is an edge.
-An attributed path is a path where each node has a list of attributes, referred to as a modifier.
+An attributed path is a path where each node has some attributes, referred to as modifiers.
 Figure 1b presents an example of a length-one attributed path, “conference papers written by Microsoft researchers”, and Figure 1c is an instance of that path.
@@
 Given a path $\mathcal{P}=t_{1}\xrightarrow{r_{1}}t_{2}\xrightarrow{r_{2}}...\xrightarrow{r_{l}}t_{l+1}$ in $\mathcal{G}$, where $l\geq 0$, a path hypothesis is defined as:
-[…]
-[…]
+$H_{path}:P^{o}_{c}\left(\text{agg}(f_{\mathcal{P}}\mid M_{t_{i}},\forall t_{i}\text{ on }\mathcal{P})\right)$
+where $o\in\{=,<>,>,<\}$; $c\in\mathbb{R}$ is a constant value; $P^{o}_{c}$ is the predicate in the format: equal, unequal, greater, or less than a value; $f_{\mathcal{P}}$ is any function of node and/or edge attributes on $\mathcal{P}$; $\text{agg}(f_{\mathcal{P}}\mid M_{t_{i}},\forall t_{i}\text{ on }\mathcal{P})$ is an aggregation function applying on the $f_{\mathcal{P}}$ of all paths whose nodes satisfy the corresponding modifiers.
 In the DBLP network defined previously, co-authorship can be represented as a path:
@@
 “The average citation of papers co-authored by Microsoft researchers is greater than 100” can be expressed as:
-[…]
+$P^{>}_{100}\left(\textit{avg}(f_{\mathcal{P}}\mid\mathtt{author}[\mathtt{MSR}],\mathtt{paper}[],\mathtt{author}[\mathtt{MSR}]\right)$
 where $f_{\mathcal{P}}=\mathtt{paper}[\mathtt{citation}]$, ${\tt MSR}$ stands for microsoft research.
-When $\mathcal{P}$ has length zero, a path hypothesis is reduced to a node hypothesis.
-An example of node hypothesis in the DBLP dataset can be “the average number of citations for conference papers is larger than 50” expressed as:
-[…]
-When $\mathcal{P}$ has length one, we refer it as an edge hypothesis.
-For instance, the edge hypothesis “the average $\mathtt{FOS\_weight}$ of conference papers on data mining is larger than 0.5” can be expressed as:
-[…]
+When $l=0$, a path hypothesis is reduced to a node hypothesis. An example in DBLP, “the average number of citations for conference papers is larger than 50”, can be expressed as:
+$P^{>}_{50}\left(\textit{avg}(\mathtt{paper}[\mathtt{citation}]\mid\mathtt{paper}[\mathtt{conference}])\right)$
+When $l=1$, we refer it as an edge hypothesis. For instance, the edge hypothesis, “the average $\mathtt{FOS\_weight}$ of conference papers on data mining is larger than 0.5”, can be expressed as:
+$P^{>}_{0.5}\left(\textit{avg}(\mathtt{WithDomain}[\mathtt{FOS\_Weight}]\mid\mathtt{paper}[\mathtt{C}],\mathtt{FOS}[\mathtt{DM}]\right)$
 C (resp. DM) stands for conference papers (resp. DM data mining).
@@
 Given an attributed graph $\mathcal{G}$, our aim is to verify a node, edge, or path hypothesis $H$ and return true or false.
-To address our problem, we formulate two objectives: O1 the hypothesis testing should be as accurate as possible; O2 the overall execution time should be as small as possible.
-When testing a hypothesis such as “the average number of citations for conference papers is larger than 50”, a common approach includes collecting representative sample data and computing an aggregate value.
-In this case, the aggregate value corresponds to the average number of citations.
-However, directly collecting all relevant, i.e., satisfying the modifiers in the hypothesis, nodes, edges, or paths from $\mathcal{G}$ and computing an aggregate value can be time-consuming.
-Therefore, we adopt the commonly used approach of sampling the graph $\mathcal{G}$ into a smaller representative graph $\mathcal{S}$.
-This assumes a sampling budget $B$ that reflects the maximum size of $\mathcal{S}$.
-For the sake of simplicity, unless stated otherwise, we assume sampling an edge or a node requires the same unitary cost $1$.
-A hypothesis estimator returns the aggregated value of nodes, edges, or paths in a hypothesis.
-The challenge of addressing O1 is that depending on the sampling method and the budget $B$, the hypothesis estimator obtained from $\mathcal{S}$ may differ from that of $\mathcal{G}$.
-We propose to first examine the applicability of existing sampling methods for graphs.
-Our observation is that the accuracy of those methods increases slowly as $B$ increases, resulting in longer times to achieve satisfactory accuracy.
-We conjecture that this is attributed to the lack of awareness of the underlying hypothesis during sampling.
-For instance, $\mathcal{S}$ should preserve as many relevant paths from $\mathcal{G}$ as possible for a path hypothesis.
-Hence, to address objective O2, we seek to design a sampling method that is aware of $H$ and preserves relevant nodes, edges, or paths in $H$.
-We further aim to demonstrate both theoretically and empirically that using our sampling, the hypothesis estimator on resulting $\mathcal{S}$ converges to its corresponding value in $\mathcal{G}$ as $B$ increases.
+We formulate two objectives to address our problem: O1 - ensure hypothesis testing is as accurate as possible; O2 - minimize execution time.
+When testing a hypothesis like “the average number of citations for conference papers is greater than 50”, a conventional approach is to collect representative conference papers from the database, compute the average number of citations, and perform the statistical test.
+However, the entire graph may not be accessible, or collecting all relevant nodes, edges, or paths from $\mathcal{G}$ is time-consuming and impractical.
+Hence, we adopt graph sampling techniques to sample a subgraph $\mathcal{S}$ from $\mathcal{G}$, assuming a sampling budget $B$ that reflects the maximum size of $\mathcal{S}$.
+For simplicity, unless stated otherwise, sampling an edge or a node incurs the same unitary cost of one.
+A hypothesis estimator computes an aggregate value for nodes, edges, or paths based on $H$.
+It directly impacts the accuracy of hypothesis testing.
+The challenge for hypothesis-agnostic samplers to achieve O1 is that they can miss relevant nodes, edges, or paths when $B$ is small.
+By examining the applicability of existing hypothesis-agnostic samplers, we observe that a large $B$ (almost the size of $\mathcal{G}$) is required to achieve an accuracy of one, which results in longer execution time and the challenge behind O2.
+We conjecture that this is due to a lack of awareness of the underlying hypothesis during sampling.
+Hence, to address two challenges, we aim to design a hypothesis-aware sampler that preserves relevant nodes, edges, or paths in $S$.
+We also intend to show, both theoretically and empirically, that using our sampler, the hypothesis estimator on $\mathcal{S}$ will converge earlier to its corresponding value in $\mathcal{G}$ as $B$ increases.
 3. Sampling-based Hypothesis Testing
@@
 After reviewing existing hypothesis-agnostic methods, we introduce our hypothesis-aware sampler, PHASE, in Section 3.2.1.
-It addresses objective O1 by ensuring convergence of hypothesis estimators, and objective O2 by incorporating hypothesis awareness into the sampling.
-Additionally, we implement two optimizations, $\text{PHASE}_{\text{opt}}$, to reduce the running time of PHASE in Section 3.2.2.
+It achieves O1 by incorporating hypothesis awareness into the sampling, and O2 by ensuring earlier convergence of hypothesis estimators.
+Additionally, we implement two optimizations in $\text{PHASE}_{\text{opt}}$ to reduce the execution time of PHASE in Section  3.2.2.
 3.1. Hypothesis-Agnostic Samplers
-Figure2 illustrates the sampling-based hypothesis testing framework we propose to test node, edge, and path hypotheses on attributed graphs.
-Based on whether the graph sampling method is hypothesis-agnostic or hypothesis-aware, there are two workflows indicated by dashed and dash-dot-dash arrows, respectively.
-Their distinction lies in where the hypothesis $H$ is introduced.
-In the hypothesis-aware workflow, $H$ is necessary during the sampling step, while in the hypothesis-agnostic workflow, $H$ is directly introduced before hypothesis testing.
-Following dashed arrows from the left, the first procedure is to sample a graph $\mathcal{S}$ using a hypothesis-agnostic sampler, which requires $\mathcal{G}$ and $B$ only.
-The existing hypothesis-agnostic samplers can be categorised into three categories: node samplers that choose $B$ nodes from $\mathcal{G}$(Stumpf et al., 2005; Adamic et al., 2001), edge samplers that choose $B$ edges from $\mathcal{G}$(Krishnamurthy et al., 2005), and random walk based samplers that pick edges and nodes through the process of random walks(Gjoka et al., 2010; Lee et al., 2012; Leskovec and Faloutsos, 2006; Li et al., 2019).
-The second procedure is to conduct the hypothesis testing using $\mathcal{S}$ and $H$.
-Specifically, we extract all relevant nodes, edges, or paths from $\mathcal{S}$ and compute the required aggregated values for hypothesis testing.
-Finally, the acceptance or rejection result, p-value, and confidence interval are returned.
+Figure 2 illustrates our sampling-based hypothesis testing framework for testing node, edge, and path hypotheses on attributed graphs.
+It consists of two steps: (1) Sampling and (2) Hypothesis Testing.
+There are two workflows: hypothesis-agnostic, shown with dashed arrows, where $H$ is considered only in the hypothesis testing step, and hypothesis-aware, indicated by dash-dot-dash arrows, which requires $H$ during the sampling step.
+In the hypothesis-agnostic workflow (dashed arrows), given $\mathcal{G}$ and $B$, a hypothesis-agnostic sampler is used in the sampling step to obtain $\mathcal{S}$.
+Existing hypothesis-agnostic samplers fall into three categories: node samplers that choose $B$ nodes from $\mathcal{G}$ (Stumpf et al., 2005; Adamic et al., 2001), edge samplers that choose $B$ edges from $\mathcal{G}$ (Krishnamurthy et al., 2005), and random walk based samplers that pick edges and nodes by random walks (Gjoka et al., 2010; Lee et al., 2012; Leskovec and Faloutsos, 2006; Li et al., 2019).
+In the hypothesis testing step, relevant nodes, edges, or paths are extracted from $\mathcal{S}$ for hypothesis testing.
+Finally, the acceptance result, p-value, and confidence interval are returned.
 3.2. Hypothesis-Aware Samplers
 3.2.1. PHASE Algorithm
-We introduce PHASE, our Path-Hypothesis-Aware SamplEr (Algorithm1).
-PHASE aims to preserve nodes, edges, or paths specified in a hypothesis $H$.
-It can be integrated with any random walk based sampler that maintains the path-preserving property.
-To ease understanding, we illustrate it with FrontierS(Ribeiro and Towsley, 2010).
-FrontierS picks a node from $m$ random seeds based on degree-proportional probability and performs a random walk by uniformly selecting a neighbor to proceed.
-It repeats the process until reaching the sampling budget $B$.
+We introduce PHASE, our Path-Hypothesis-Aware SamplEr (Algorithm 1), for the sampling step.
+Since $\mathcal{G}$ may contain both relevant and irrelevant nodes, edges, or paths for a given $H$, PHASE aims to preferentially sample those specified in $H$.
+This strategy resembles stratified sampling (Cochran, 1977; Neyman, 1992), where the target population is sampled at higher rates without bias.
+PHASE can be integrated with any random walk based sampler.
+To clarify, we describe it using Frontier Sampler (FrontierS(Ribeiro and Towsley, 2010), which picks a node from $m$ random seeds based on degree-proportional probability and performs a random walk by uniformly selecting a neighboring node.
+This process repeats until $B$ is reached.
 Algorithm 1 PHASE
-1:Input: $\mathcal{G}=(\mathcal{V,E})$, $B$, $H$
+1:Input: $\mathcal{G}=(\mathcal{V,E})$, $B$, $H$, $Q$
 2:Output: a sampled graph $\mathcal{S}$
 3:Initialize $L=(v_{1},v_{2},\ldots,v_{m})$ with $m$ randomly chosen nodes
-4:$L_{w}=\text{AssignWeight}(L,H)$triangleright Assign $w_{h}$ to $x_{1}$ nodes and $w_{l}$ to other nodes.
+4:$L_{w}=\text{AssignWeight}(L,H)$triangleright Assign $w_{h}$ to $x_{1}$ nodes and $w_{l}$ to others.
 5:$\mathcal{V}_{\mathcal{S}}=\{\}$, $\mathcal{E}_{\mathcal{S}}=\{\}$
@@
 9:    $N\leftarrow N[v]$triangleright $N[v]$ is the set of neighbors of $v$.
-10:    $N_{w}=\text{AssignWeight}(N,H)$triangleright Assign weights to nodes in $N$ according to the transition probability matrices.
-11:    Select $u\in N$ with probability $N_{w}$
+10:    $N_{w}=\text{AssignWeight}(N,H,{Q})$triangleright Assign weights according to $Q$.
+11:    Select $u\in N$ with the normalized $N_{w}$
 12:    $\mathcal{V}_{\mathcal{S}}$.update($v,u$)
@@
 16:return $\mathcal{S}=\{\mathcal{V}_{\mathcal{S}},\mathcal{E}_{\mathcal{S}}\}$
-PHASE takes $\mathcal{G}$, $B$, and $H$ as inputs.
+PHASE takes $\mathcal{G}$, $B$, $H$, $Q$ as inputs.
 For a node, edge, or path hypothesis, $H$ contains $\mathcal{P}$ with lengths zero, one, or more than one, respectively.
-It initializes $m$ seed nodes randomly in a list $L$ (Line 1), which increases the probability of selecting relevant nodes, edges, and paths from $\mathcal{G}$.
+$m$ seed nodes in $\mathcal{G}$ are randomly initialized and stored in a list $L$ (Line 1).
+This line increases the chance of picking relevant nodes, edges, and paths and ensures they are not locally clustered, preventing locality bias.
 Each seed is assigned a weight in $L_{w}$ (Line 2) to guide the selection of seed nodes for random walks later.
-These weights are determined based on heuristics, with nodes satisfying the first node modifier on $\mathcal{P}$ (denoted by $x_{1}$) assigned a higher weight $w_{h}$ and all other nodes assigned a lower weight $w_{l}$.
+The weights are determined based on heuristics: nodes satisfying the first node modifier on $\mathcal{P}$ (denoted by $x_{1}$) receive a higher weight $w_{h}$, while others receive a lower weight $w_{l}$, where $w_{h}\geq w_{l}>0$.
 Lines 4-14 describe an iterative random walk to select nodes until $B$ is reached.
-During each iteration, the algorithm chooses a node $v$ from $L$ with probability $L_{w}$ (Line 6).
-Then, it selects one of $v$’s neighbors $u$ using a weighted selection process (Lines 7-9).
-Specifically, in Line 9, we adjust the transition probability to bias the random walks towards nodes, edges, or paths that satisfy $\mathcal{P}$ in $H$.
-The transition probability matrices for node, edge, and path hypotheses are depicted in Figures 3a, 3b, and 3c, respectively.
+During each iteration, the algorithm chooses a node $v$ from $L$ based on the probability distribution $L_{w}$ (Line 6).
+Next, it picks a neighbor $u$ of $v$ using a weighted selection process (Lines 7-9) (O1).
+Specifically, in Line 8, we steer the random walks towards relevant nodes, edges, or paths.
+The transition probability matrices for node, edge, and path hypotheses are shown in Figures  3a,  3b, and  3c, respectively.
+For a node hypothesis, a higher weight $w_{h}$ is assigned to $x_{1}$ regardless of the current node.
+For an edge hypothesis, $w_{h}$ is given to $x_{2}$ when the current node is $x_{1}$, and given to $x_{1}$ in other cases.
+This increases the chance of sampling more relevant edges $(x_{1},x_{2})$.
+The random walks are 1st-order in Figures  3a and  3b.
+Path hypotheses ($l=2$) have 2nd-order random walks (Figure  3c), meaning the transition depends on the current and previous nodes.
+For example, given two nodes $x_{1}$ and $x_{2}$, $x_{3}$ gets a higher weight $w_{h}$.
+The consistent $w_{h}$ prevents sampling bias within the population.
 After adding nodes $u$ and $v$ to $\mathcal{V}_{S}$ (Line 10), the algorithm updates $L$ by replacing $v$ with $u$ (Lines 11) and decreases the sampling budget by one (Lines 12).
@@
 The resulting sampled graph $\mathcal{S}$ is the induced subgraph from $\mathcal{V}_{S}$.
-$x_{1}$$y$$x_{1}$$w_{h}$$w_{l}$$y$$w_{h}$$w_{l}$(a)$x_{1}$$x_{2}$$y$$x_{1}$$w_{l}$$w_{h}$$w_{l}$$x_{2}$$w_{h}$$w_{l}$$w_{l}$$y$$w_{h}$$w_{l}$$w_{l}$(b)$x_{1}$$x_{2}$$x_{3}$$y$$x_{1},x_{1}$$w_{l}$$w_{h}$$w_{l}$$w_{l}$$x_{1},x_{2}$$w_{l}$$w_{l}$$w_{h}$$w_{l}$$x_{1},x_{3}$$w_{h}$$w_{l}$$w_{l}$$w_{l}$$x_{1},y$$w_{h}$$w_{l}$$w_{l}$$w_{l}$$\vdots$(c)
+$x_{1}$$y$$x_{1}$$w_{h}$$w_{l}$$y$$w_{h}$$w_{l}$(a)$x_{1}$$x_{2}$$y$$x_{1}$$w_{l}$$w_{h}$$w_{l}$$x_{2}$$w_{h}$$w_{l}$$w_{l}$$y$$w_{h}$$w_{l}$$w_{l}$(b)$x_{1}$$x_{2}$$x_{3}$$y$$x_{1},x_{1}$$w_{l}$$w_{h}$$w_{l}$$w_{l}$$x_{1},x_{2}$$w_{l}$$w_{l}$$w_{h}$$w_{l}$$x_{1},x_{3}$$w_{h}$$w_{l}$$w_{l}$$w_{l}$$x_{1},y$$w_{h}$$w_{l}$$w_{l}$$w_{l}$(c)
 Figure 3.
-Transition probability matrices for node (a), edge (b), and path ($l=2$) (c) hypotheses, where $w_{h}\geq w_{l}>0$ denote transition probabilities.
-$x_{i}$ represents nodes in $\mathcal{G}$ satisfying the $i$-th node modifier on $\mathcal{P}$, while $y$ represents nodes not satisfying any node modifier on $\mathcal{P}$.
-(a) and (b) involve 1st-order random walks, whereas (c) involves 2nd-order random walks because the probability of selecting the next node depends on both the current and previous nodes.
+Transition probability matrices ${Q}$ for (a) node, (b) edge, and (c) path $(l=2)$ hypotheses (up to the first four rows).
+$x_{i}$ represents nodes in $\mathcal{G}$ satisfying the $i$-th node modifier on $\mathcal{P}$ and $y$ represents other nodes.
 3.2.2. $\text{PHASE}_{\text{opt}}$ Algorithm
-We further propose optimizations to reduce execution time.
+We further propose optimizations to reduce execution time based on heuristics.
 Algorithm 2 introduces two lines that replace line 7 in Algorithm 1.
-Firstly, in densely connected graphs, the computation of neighbor weights in step 8 of Algorithm 1 can be time-consuming due to the potentially large number of neighbors.
-To mitigate this, we randomly sample a subset of $\min\{|N|,n\}$ neighbors as candidates for selection, where $n$ is a parameter.
-Secondly, we adopt a non-backtracking approach by avoiding the selection of visited nodes during the sampling process(Lee et al., 2012).
-This strategy prevents cycles and unnecessary node visits.
+In line 1, we employ a non-backtracking approach to avoid selecting previously visited nodes during sampling (Lee et al., 2012), preventing cycles and unnecessary revisits.
+Additionally, in densely connected graphs, computing neighbor weights in line 8 of Algorithm 1 can be time-consuming due to the potentially large number of neighbors.
+To address this, line 2 in Algorithm 2 introduces random sampling of a subset of $\min\{|N^{\prime}|,n\}$ neighbors as candidates, where $n$ is a parameter.
+This can effectively reduce computation time.
+But it may result in some accuracy loss when $n$ is significantly smaller than $|N^{\prime}|$.
 Algorithm 2 $\text{PHASE}_{\text{opt}}$
-(showing only 2 lines that replace line 7 in Algorithm 1
-1:$N^{\prime}\leftarrow N[v]-\mathcal{V_{S}}$ (Optim 2)
-2:$N\leftarrow$ Select $\text{min}\{|N^{\prime}|,n\}\text{ nodes randomly from }N^{\prime}$ (Optim 1)
+(showing only 2 lines that replace line 7 in Algorithm  1
+1:$N^{\prime}\leftarrow N[v]-\mathcal{V_{S}}$ (Optim 1)
+2:$N\leftarrow$ Select $\text{min}\{|N^{\prime}|,n\}\text{ nodes randomly from }N^{\prime}$ (Optim 2)
 Complexity.
@@
 3.2.3. Convergence of Hypothesis Estimators
-Let us demonstrate the convergence of hypothesis estimators using our proposed sampling algorithms (Challenge 1).
-By Lemma 5.1 from(Ribeiro and Towsley, 2010), FrontierS is equivalent to the sampling process of a SRW over $\mathcal{G}^{m}=(\mathcal{V}^{m},\mathcal{E}_{m})$, where $\mathcal{G}^{m}$ is the $m$-th Cartesian power of $\mathcal{G}$ and
-$\mathcal{V}^{m}=\{(v_{1},\ldots,v_{m})|v_{1}\in\mathcal{V}\wedge\dots\wedge v_{m}\in\mathcal{V}\}$
-is the $m$-th Cartesian power of $\mathcal{V}$.
-$\forall\boldsymbol{u,v}\in\mathcal{V}^{m}$, $(\boldsymbol{u,v})\in\mathcal{E}_{m}$ if there exists $i$ such that $(u_{i},v_{i})\in\mathcal{E}$ and for all $j\neq i$, $u_{j}=v_{j}$.
-Given $\mathcal{G}$ is a directed, connected, and non-bipartite graph, in the steady state, the sampled edges of FrontierS form a stationary sequence and satisfy the following Strong Law of Large Numbers (SLLN), the proof of which can be found in (Ribeiro and Towsley, 2010).
+We demonstrate the convergence of hypothesis estimators, which ensures the convergence of hypothesis testing accuracy to one, for all sampling methods.
+Then, we show our proposed sampler achieves earlier and smoother convergence (O2).
+We will focus on scenarios where agg is an average function to construct and prove the convergence of hypothesis estimators.
+Estimators for other aggregation functions, such as maximum and minimum, can be derived analogously.
+For a path hypothesis ($l\geq 0$), $\text{avg}(f_{\mathcal{P}}\mid M_{t_{i}},\forall t_{i}\text{ on }\mathcal{P})$ has a primary subject $f_{\mathcal{P}}$.
+Let $\mathcal{P}^{*}$ be all relevant paths in $\mathcal{G}$, and $\mathcal{P_{G}}$ be all paths with the same length as $\mathcal{P}$ in $\mathcal{G}$.
+The mean value of the path hypothesis, $\theta_{path}$, is
+(1) $\theta_{path}=\frac{1}{|\mathcal{P}^{*}|}\sum_{\forall\mathcal{P}\in\mathcal{P_{G}}}T(\mathcal{P})$
+where $T(\mathcal{P})=f_{\mathcal{P}}\times\mathbbm{1}_{M_{t_{i}}\subseteq\mathcal{L}_{\phi(t_{i})}\forall t_{i}\text{ on }\mathcal{P}}$.
+Replacing $\mathcal{G}$ with the sampled graph $\mathcal{S}$, the estimator for $\theta_{path}$ on $\mathcal{S}$ is
+(2) $\hat{\theta}_{path}=\frac{1}{|\mathcal{P}^{*}\cap\mathcal{P_{S}}|}\sum_{\forall\mathcal{P}\in\mathcal{P_{S}}}T(\mathcal{P})$
+When $l=0$ (resp. $l=1$), we name the corresponding mean value $\theta_{node}$ (resp. $\theta_{edge}$) and estimator $\hat{\theta}_{node}$ (resp. $\hat{\theta}_{edge}$).
+Some sampling methods, including random edge sampler, simple random walk, and FrontierS, obey the Strong Law of Large Numbers (SLLN)  (Ribeiro and Towsley, 2010).
+Their $\hat{\theta}_{node}$ and $\hat{\theta}_{edge}$ are asymptotically unbiased estimators of $\theta_{node}$ and $\theta_{edge}$, respectively, when $B\to\infty$.
 Theorem 1 (SLLN).
@@
 almost surely, i.e. the event occurs with probability one.
-For a path hypothesis with length $l$, it involves $(l-1)$-th order random walk, which is equivalent to a SRW over $\mathcal{G}^{l-1}=(\mathcal{V}^{l-1},\mathcal{E}_{l-1})$(Lovász, 1993).
-Hence, PHASE’s random walks can be viewed as a SRW over $\mathcal{G}^{l-1+m}=(\mathcal{V}^{l-1+m},\mathcal{E}_{l-1+m})$.
-As PHASE ensures that the resulting random walks over $\mathcal{G}^{l-1+m}$ maintain both irreducibility and aperiodicity, it preserve stationarity and adhere to the SLLN.
-$\text{PHASE}_{\text{opt}}$ also follows the SLLN because the stationary distribution of the NBRW remains the same as that of the SRW.
-By Theorem1, we can construct hypothesis estimators for node, edge, and path hypotheses and prove their convergence.
-Here, we focus on scenarios where agg represents an average function.
-The hypothesis estimators for other aggregation functions, such as maximum and minimum, can be derived in a similar manner.
-A path hypothesis ($l\geq 2$): $\text{avg}(f_{\mathcal{P}}|M_{t_{i}},\forall t_{i}\text{ on }\mathcal{P})$ has a primary subject $f_{\mathcal{P}}$.
-Let $\mathcal{P}^{*}$ be all relevant paths in $\mathcal{G}$, and $\mathcal{P_{G}}$ be all paths with the same length as $\mathcal{P}$ in $\mathcal{G}$.
-The mean value of the path hypothesis, $\theta_{path}$, is
-(1) $\theta_{path}=\frac{1}{|\mathcal{P}^{*}|}\sum_{\forall\mathcal{P}\in\mathcal{P_{G}}}T(\mathcal{P})$
-where $T(\mathcal{P})=f_{\mathcal{P}}\times\mathbbm{1}_{M_{t_{i}}\subseteq\mathcal{L}_{\phi(t_{i})}\forall t_{i}\text{ on }\mathcal{P}}$.
-Replacing $\mathcal{G}$ with the sampled graph $\mathcal{S}$, the estimator for $\theta_{path}$ is
-(2) $\hat{\theta}_{path}=\frac{1}{|\mathcal{P}^{*}\cap\mathcal{P_{S}}|}\sum_{\forall\mathcal{P}\in\mathcal{P_{S}}}T(\mathcal{P})$
-When $B$ goes to infinity, $\mathcal{S}$ converges to $\mathcal{G}$. By Theorem1, we have
-(3) $\lim_{B\to\infty}\frac{1}{|\mathcal{P}^{*}\cap\mathcal{P_{S}}|}\sum_{\forall\mathcal{P}\in\mathcal{P_{S}}}T(\mathcal{P})\to\frac{1}{|\mathcal{P}^{*}|}\sum_{\forall\mathcal{P}\in\mathcal{P_{G}}}T(\mathcal{P})$
-almost surely. Hence, $\hat{\theta}_{path}$ is an asymptotically unbiased estimator of $\theta_{path}$. Hypothesis estimators for node and edge hypotheses can be derived similarly.
+In general scenarios with any path hypotheses ($l\geq 0$) and sampler, SLLN may not apply, and the estimator may not be asymptotically unbiased.
+However, $\hat{\theta}_{\text{path}}$ still converges to $\theta_{\text{path}}$ due to Finite Population Correction (FPC), a statistical adjustment made when sampling without replacement from a finite population (Bondy and Zlot, 1976).
+As $B$ increases, $\text{FPC}=\frac{|\mathcal{P}^{*}|-\mathcal{P_{S}}}{|\mathcal{P}^{*}|-1}$ approaches zero, which ensures the convergence of the hypothesis estimator to its ground truth.
+While all samplers ensure the convergence of the hypothesis estimator, our proposed samplers achieve earlier convergence than other methods.
+This is due to their higher probability of selecting relevant nodes, edges, and paths.
+For instance, when $l=2$, the probability to pick nodes $(x_{1},x_{2},x_{3})$ is:
+$P(x_{1},x_{2},x_{3})$ $=P(x_{1})\times(P(x_{2}\mid x_{1})\times A(x_{2}\mid x_{1}))$
+×$\quad\times(P(x_{3}\mid x_{1},x_{2})\times A(x_{3}\mid x_{1},x_{2}))$
+where $A(x_{i}\mid x_{i-1})$ indicates the accessibility (0 or 1) of node $x_{i}$ from $x_{i-1}$.
+Assuming node $x_{i}$ has $n_{i}$ neighbors with $k_{i}$ of type $x_{i+1}$ and there are $d$ $x_{1}$ in $m$ seed nodes, where $d\geq 0$ and $k_{i}\geq 0$:
+$P_{\text{PHASE}}(x_{1},x_{2},x_{3})$ $=\frac{d\cdot w_{h}}{d\cdot w_{h}+(m-d)\cdot w_{l}}$
+×$\quad\times\left(\frac{w_{h}}{k_{1}\cdot w_{h}+(n_{1}-k_{1})\cdot w_{l}}\times A(x_{2}\mid x_{1})\right)$
+×$\quad\times\left(\frac{w_{h}}{k_{2}\cdot w_{h}+(n_{2}-k_{2})\cdot w_{l}}\times A(x_{3}\mid x_{1},x_{2})\right)$
+Random node and edge samplers have the lowest probability of picking $(x_{1},x_{2},x_{3})$.
+Random walk based samplers have a transitional probability from $x_{i}$ of $\frac{1}{n_{i}}\leq\frac{w_{h}}{k_{i}\cdot w_{h}+(n_{i}-k_{i})\cdot w_{l}}$, particularly when $k_{i}$ is large.
+With a higher probability to pick $(x_{1},x_{2},x_{3})$, PHASE and $\text{PHASE}_{\text{opt}}$ have earlier convergence of the hypothesis estimator than existing hypothesis-agnostic samplers.
+Also, as shown in Figure 4, the empirical results of two path hypotheses, whose details are in Table 2, align with the claim.
+Figure 4. The convergence of hypothesis estimator for two path hypotheses: DB-P3 (left) and YP-P3 (right).
 4. Experiments
-The goal of our experiments is twofold: 1) test the effectiveness of optimizations in $\text{PHASE}_{\text{opt}}$ compared to PHASE (Section 4.3), and 2) compare hypothesis-agnostic and hypothesis-aware samplers in terms of test significance (Section 4.4), accuracy (Section 4.5), and time (Section 4.6).
-The source code is publicly available  [^1] https://github.com/Carrieww/GraphHT.
+The goal of our experiments is twofold: 1) test the effectiveness of optimizations in $\text{PHASE}_{\text{opt}}$ compared to PHASE (Section  4.3), and 2) compare hypothesis-agnostic and hypothesis-aware samplers in terms of test significance (Section  4.4), accuracy (Section  4.5), and time (Section  4.6).
 4.1. Experimental Setup
@@
 Dataset #(Nodes) #(Edges) Density #(Node Types) #(Edge Types)
-MovieLens 9,705 996,656 1.06e-02
-2
-1
-DBLP 1,623,013 11,040,170 4.19e-06
-4
-4
-Yelp 2,136,118 6,743,879 1.48e-06
-2
-1
-Datasets. We used three datasets (Harper and Konstan, 2016; Tang et al., 2008), extracted from real attributed networks: MovieLens, DBLP, and Yelp. Table 1 shows the statistics of the datasets.
-Samplers We compare $\text{PHASE}_{\text{opt}}$ with eleven existing samplers:
-  • Node samplers: Random Node Sampler (RNS), Degree-Based Sampler (DBS)
-  • Edge samplers: Random Edge Sampler (RES)
-  • Random walk based samplers: Simple Random Walk (SRW), Frontier Sampler (FrontierS), Non-Backtracking Random Walk (NBRW), Random Walk with Restarter (RWR), Metropolis-Hastings Random Walk (MHRW), Snaw Ball Sampler (SBS), Forest Fire Sampler (FFS), Shortest Path Sampler (ShortestPathS)
+MovieLens 9,705 996,656 1.06e-02 2 1
+DBLP 1,623,013 11,040,170 4.19e-06 4 4
+Yelp 2,136,118 6,743,879 1.48e-06 2 1
+Datasets.
+We use three datasets (Harper and Konstan, 2016; Tang et al., 2008; yel, 2015), extracted from real attributed networks: MovieLens, DBLP, and Yelp.
+Table 1 shows the statistics of the datasets.
+Samplers We compare $\text{PHASE}_{\text{opt}}$ with 11 existing samplers:
+  • Node samplers: Random Node Sampler (RNS(Stumpf et al., 2005), Degree-Based Sampler (DBS)  (Adamic et al., 2001)
+  • Edge samplers: Random Edge Sampler (RES)  (Krishnamurthy et al., 2005)
+  • Random walk based samplers: Simple Random Walk (SRW)  (Gjoka et al., 2010), Frontier Sampler (FrontierS)  (Ribeiro and Towsley, 2010), Non-Backtracking Random Walk (NBRW)  (Lee et al., 2012), Random Walk with Restarter (RWR)  (Leskovec and Faloutsos, 2006), Metropolis-Hastings Random Walk (MHRW)  (Stutzbach et al., 2009), Snaw Ball Sampler (SBS)  (Goodman, 1961), Forest Fire Sampler (FFS)  (Leskovec and Faloutsos, 2006), Shortest Path Sampler (ShortestPathS)  (Rezvanian and Meybodi, 2015)
 Hypotheses.
-For each dataset (DB) and hypothesis type (N for node hypotheses), we provide three example hypotheses (labeled as DB-N1) in experiments.
-For example, on Yelp, we explore an edge hypothesis testing if “fast food average ratings exceed 4”.
-In DBLP, we examine a path hypothesis to see if “papers co-authored by Chinese institutes on data mining have over 50 citations”.
-To assess path length effects, we further test two path hypotheses with lengths of three and four in DBLP.
+For each dataset (e.g., DB) and hypothesis type (e.g., N for node hypotheses), we chose three example hypotheses (e.g., DB-N1) in the experiment.
+Due to the page limit, examples for DBLP and Yelp are shown in Table 2.
+These hypotheses are chosen based on context and difficulty.
+Based on the context of the datasets, the subject of interest and the constant in the hypothesis should reflect meaningful requests from real users.
+The difficulty is related to the path length and the number of relevant nodes, edges, or paths, as indicated in the third column of Table 2.
+The longer the path or the fewer the relevant nodes, edges, or paths, the more difficult it becomes to sample them from $\mathcal{G}$ for accurate hypothesis testing.
+We use the following path hypotheses of DBLP with lengths of three and four in the experiment: “the average citation of two papers, each authored by a Microsoft researcher and citing each other, $>$ 50” and “the average citation of two conference papers, each authored by a Microsoft researcher, $>$ 50”.
+Table 2. Examples of hypotheses for DBLP and Yelp
+Hypothesis
+Type
+Example Relevant nodes,
+edges, paths
+DBLP
+Node DB-N1: The avg citation of papers published as journals $>$ 20 199205 (easy)
+DB-N2: The avg citation of papers in conferences in 2010 $>$ 10 31566 (medium)
+DB-N3: The avg citation of papers published in Journal in 2017 $>$ 10 1588 (hard)
+Edge DB-E1: The avg weight of conference papers on data mining $>$ 0.5 44925 (easy)
+DB-E2: The avg weight of journal papers on data mining $>$ 0.5 13400 (medium)
+DB-E3: The avg weight of papers on telecommunications network $>$ 0.5 2510 (hard)
+Path DB-P1: The avg weight of papers by China’s institutes on data mining $>$ 0.5 17671 (easy)
+DB-P2: The avg citation of papers co-authored by authors in Peking and China’s institutions $>$ 50 7065 (medium)
+DB-P3: The avg citation of conference papers by Microsoft Researchers $>$ 50 3217 (hard)
+Yelp
+Node YP-N1: The avg reviews given by users with high popularity $>$ 200 112043 (easy)
+YP-N2: The avg stars given by users who have low prolificacy and medium popularity $>$ 3 16429 (medium)
+YP-N3: The avg number of reviews of business in Illinois $>$ 4 2144 (hard)
+Edge YP-E1: The avg ratings of fast food $>$ 4 224536 (easy)
+YP-E2: The avg ratings of furniture stores $>$ 3 33040 (medium)
+YP-E3: The avg percentage of useful reviews given by useful writers to Illinois businesses $>$ 0.5 4242 (hard)
+Path YP-P1: The avg rating difference on path [business in FL - high popularity user - business in LA] $>$ 0.5 615174 (easy)
+YP-P2: The avg rating difference on path [business in LA - high popularity user - business in IL] $>$ 0.5 15542 (medium)
+YP-P3: The avg rating difference on path [business in LA - medium popularity user - business in AB] $>$ 0.5 1080 (hard)
 Parameter Choice.
-For PHASE and $\text{PHASE}_{\text{opt}}$, We set $m=50$ to maintain the path-preserving capability of multi-dimensional random walks, and $n=30$ to strike a balance between accuracy and time efficiency.
-Additionally, we choose $w_{h}=10$ and $w_{l}=0.1$ to regulate the bias towards relevant nodes, edges, and paths.
+We determine optimal settings for $m$, $n$, $w_{h}$, and $w_{l}$ through a grid search, with $m$ and $n$ ranging from 10 to 200, and $w_{h}$ and $w_{l}$ from 0.1 to 20.
+We set $m=50$ to maintain the path-preserving ability of multi-dimensional random walks, and $n=30$ to strike a balance between accuracy and time efficiency.
+$w_{h}=10$ and $w_{l}=0.1$ help regulate the prioritization towards sampling relevant nodes, edges, or paths.
 The sampling budget $B$ is maintained as a proportion of the total number of nodes in $\mathcal{G}$ for all samplers.
+In real deployment, we recommend iteratively increasing $B$ until the accuracy stabilizes at a high threshold (e.g., 0.9) based on the average from 30 samples to determine the optimal $B$.
 As for existing hypothesis-agnostic samplers, we use the best parameters in their respective settings.
@@
 4.2. Evaluation Measures
-Accuracy.
-Accuracy at a specific sampling proportion reflects the effectiveness of a sampling method.
-The accuracy below measures the number of matched hypothesis testing results on $\mathcal{G}$ and $\mathcal{S}$:
+Accuracy. Accuracy at a specific sampling proportion reflects the effectiveness of a sampling method. It measures the number of matched hypothesis testing results on $\mathcal{G}$ and $\mathcal{S}$:
 $\text{Accuracy}=\frac{1}{k}\sum_{k}\mathbbm{1}_{H(\mathcal{G})==H(\mathcal{S})}$
 where $H(\mathcal{G})$ and $H(\mathcal{S})$ return 0 (false) or 1 (true).
-Time. We measure the total execution time, including both the sampling time and the time taken to extract relevant information from $\mathcal{S}$ for hypothesis testing.
-Table 2. Average execution time (sec) of PHASE and $\text{PHASE}_{\text{opt}}$ for DBLP node, edge, and path hypotheses.
-Node Edge Path
-PHASE 115.66 539.15 441.3
-$\text{PHASE}_{\text{opt}}$ 5.56 8.76 5.91
-Table 3. The accuracy of eleven existing samplers and $\text{PHASE}_{\text{opt}}$ on three datasets and three types of hypothesis.
-Dataset
-Hypothesis Type
-Sampling Proportion (%)
-$\text{PHASE}_{\text{opt}}$ RES RNS DBS SRW NBRW RWR MHRW ShortestPathS FrontierS FFS SBS
-Node
-1
-0.9 0.89 0.83 0.86 0.88 0.89 0.87 0.89 0.86 0.84 0.83 0.56
-Edge
-2.5
-0.98 0.68 0.77 0.98 0.99 1 0.99 0.98 0.99 0.73 0.93 0.91
-MovieLens
-Path
-5
-0.99 0.1 0.88 0.83 0.82 0.89 0.91 0.98 0.95 0.38 0.85 0.62
-Node
-0.2
-0.96 0.48 0.87 0.93 0.92 0.91 0.9 0.94 0.92 0.92 0.94 0.88
-Edge
-0.2
-0.76 0.48 0 0.71 0.73 0.69 0.7 0.29 0.69 0.63 0.7 0.42
-DBLP
-Path
-0.2
-0.89 0 0 0 0.26 0.3 0.32 0.18 0.33 0.043 0.3 0.12
-Node
-0.1
-0.99 0.65 0.77 0.61 0.69 0.69 0.69 0.77 0.7 0.64 0.66 0.48
-Edge
-1
-1 0.73 0.54 0.76 0.79 0.91 0.87 0.84 0.76 0.77 0.79 0.71
-Yelp
-Path
-1
-0.99 0.11 0.05 0.79 0.78 0.78 0.72 0.8 0.99 0.42 0.67 0.42
-Table 4. The execution time (sec) of eleven existing samplers and $\text{PHASE}_{\text{opt}}$ on three datasets and three types of hypothesis.
-Dataset
-Hypothesis Type
-Sampling Proportion (%)
-$\text{PHASE}_{\text{opt}}$ RES RNS DBS SRW NBRW RWR MHRWS ShortestPathS FrontierS FFS SBS
-MovieLens
-Node
-1
-0.083 0.99 0.023 0.077 0.057 0.06 0.05 0.06 0.063 0.083 0.067 0.047
-Edge
-2.5
-0.45 0.99 0.11 0.36 0.31 0.36 0.33 0.29 0.26 0.33 0.35 0.31
-Path
-5
-4.92 1.03 0.34 4.32 4.53 4.55 4.76 3.10 0.95 0.38 3.87 3.08
-DBLP
-Node
-0.2
-5.56 18.70 0.55 6.98 7.98 8.22 9.67 1.66 31.32 10.48 5.73 3.03
-Edge
-0.2
-8.76 22.57 0.90 10.07 14.76 12.89 12.32 3.53 33.91 14.46 8.87 5.27
-Path
-0.2
-5.44 19.61 0.71 6.87 8.24 8.39 8.13 2.17 31.13 9.37 5.33 3.01
-Yelp
-Node
-0.1
-2.56 13.84 1.02 6.81 1.19 0.96 1.16 1.12 1.84 1.68 1.02 6.58
-Edge
-1
-19.42 16.88 2.01 13.61 8.58 8.97 9.17 6.93 30.23 10.22 10.27 6.03
-Path
-1
-56.97 14.35 1.53 23.48 15.97 16.61 19.03 7.35 37.81 9.95 25.61 19.94
+Time. We measure the total execution time, including the sampling time, the time taken to extract relevant information from $\mathcal{S}$ for hypothesis testing, and hypothesis testing time.
+Table 3. The accuracy of 11 existing samplers and $\text{PHASE}_{\text{opt}}$ on three datasets and three types of hypothesis.
+Dataset Hypothesis Type Sampling Proportion (%) $\text{PHASE}_{\text{opt}}$ PHASE RES RNS DBS SRW NBRW RWR MHRW ShortestPathS FrontierS FFS SBS
+Node 1 0.9 1 0.89 0.83 0.86 0.88 0.89 0.87 0.89 0.86 0.84 0.83 0.56
+Edge 2.5 0.98 1 0.68 0.77 0.98 0.99 1 0.99 0.98 0.99 0.73 0.93 0.91
+MovieLens Path 5 0.99 1 0.1 0.88 0.83 0.82 0.89 0.91 0.98 0.95 0.38 0.85 0.62
+Node 0.2 0.96 1 0.48 0.87 0.93 0.92 0.91 0.9 0.94 0.92 0.92 0.94 0.88
+Edge 0.2 0.76 0.79 0.48 0 0.71 0.73 0.69 0.7 0.29 0.69 0.63 0.7 0.42
+DBLP Path 0.2 0.89 0.88 0 0 0 0.26 0.3 0.32 0.18 0.33 0.043 0.3 0.12
+Node 0.1 0.99 1 0.65 0.77 0.61 0.69 0.69 0.69 0.77 0.7 0.64 0.66 0.48
+Edge 1 1 1 0.73 0.54 0.76 0.79 0.91 0.87 0.84 0.76 0.77 0.79 0.71
+Yelp Path 1 0.99 1 0.11 0.05 0.79 0.78 0.78 0.72 0.8 0.99 0.42 0.67 0.42
+Table 4. The execution time (sec) of 11 existing samplers and $\text{PHASE}_{\text{opt}}$ on three datasets and three types of hypothesis.
+Dataset Hypothesis Type Sampling Proportion (%) $\text{PHASE}_{\text{opt}}$ PHASE RES RNS DBS SRW NBRW RWR MHRWS ShortestPathS FrontierS FFS SBS
+MovieLens Node 1 0.083 0.41 0.99 0.023 0.077 0.057 0.06 0.05 0.06 0.063 0.083 0.067 0.047
+Edge 2.5 0.45 2.07 0.99 0.11 0.36 0.31 0.36 0.33 0.29 0.26 0.33 0.35 0.31
+Path 5 4.92 14.80 1.03 0.34 4.32 4.53 4.55 4.76 3.10 0.95 0.38 3.87 3.08
+DBLP Node 0.2 5.56 418.07 18.70 0.55 6.98 7.98 8.22 9.67 1.66 31.32 10.48 5.73 3.03
+Edge 0.2 8.76 414.47 22.57 0.90 10.07 14.76 12.89 12.32 3.53 33.91 14.46 8.87 5.27
+Path 0.2 5.44 236.82 19.61 0.71 6.87 8.24 8.39 8.13 2.17 31.13 9.37 5.33 3.01
+Yelp Node 0.1 2.56 6.10 13.84 1.02 6.81 1.19 0.96 1.16 1.12 1.84 1.68 1.02 6.58
+Edge 1 19.42 67.55 16.88 2.01 13.61 8.58 8.97 9.17 6.93 30.23 10.22 10.27 6.03
+Path 1 56.97 130.3 14.35 1.53 23.48 15.97 16.61 19.03 7.35 37.81 9.95 25.61 19.94
 4.3. PHASE vs $\text{PHASE}_{\text{opt}}$
-We evaluate the effectiveness of optimizations in $\text{PHASE}_{\text{opt}}$ compared to PHASE using the DBLP dataset, as shown in Table 2.
-$\text{PHASE}_{\text{opt}}$ is at least 20 times faster than PHASE while maintaining comparable accuracy (less than 5% accuracy reduction).
-Consequently, we exclusively use $\text{PHASE}_{\text{opt}}$ for subsequent experiments.
+We evaluate $\text{PHASE}_{\text{opt}}$’s optimizations compared to PHASE in Table 3 and 4.
+For DBLP, $\text{PHASE}_{\text{opt}}$ is at least 43 times faster than PHASE with minimal accuracy loss (under 4%).
+Consequently, we use $\text{PHASE}_{\text{opt}}$ exclusively in further experiments and discussions.
 4.4. Significance
@@
 Figure 5. DBLP p-value and CI plot for the hypothesis DB-P1
-To examine statistical significance and estimation precision, we analyze the trend of p-values and confidence intervals (CI) with increasing $B$.
-We illustrate this trend using a path hypothesis on DBLP as all other hypotheses show similar trends.
-As shown in Figure 3(a), the p-value of $\text{PHASE}_{\text{opt}}$ consistently stays below the significance level (e.g., 0.05) and decreases gradually.
-This trend indicates stronger evidence against the null hypothesis, showing higher statistical significance as $B$ increases.
-Figure 3(b) shows narrowing CI with increasing $B$, indicating more precise and reliable estimations of the hypothesis estimator.
-$\text{PHASE}_{\text{opt}}$ consistently achieves the narrowest bound at any $B$, indicating the most precise and reliable estimate among all samplers.
+To assess statistical significance and estimation precision, we evaluate p-value trends and confidence intervals (CIs) as $B$ increases using DB-P3, as shown in Figure  5.
+Similar trends are observed for other hypotheses.
+$\text{PHASE}_{\text{opt}}$ consistently maintains p-values below the significance level (e.g., 0.05), showing stronger evidence against the null hypothesis as $B$ increases.
+It also exhibits the narrowest CI at all $B$, reflecting the highest precision among other samplers.
 (a) Accuracy $(l=3)$
 (b) Accuracy $(l=4)$
-Figure 5. Accuracy versus sampling proportion plots for DBLP path hypotheses with length three (a) and four (b).
+(c) Time-Accuracy $(l=3)$
+(d) Time-Accuracy $(l=4)$
+Figure 6. Accuracy (a, b) and time-accuracy (c, d) performance for DBLP path hypotheses.
+(a) DB-N1 (easy)
+(b) DB-N2 (medium)
+(c) DB-N3 (hard)
+(d) DB-N3 (hard)
+(e) DB-E1 (easy)
+(f) DB-E2 (medium)
+(g) DB-E3 (hard)
+(h) DB-E3 (hard)
+(i) DB-P1 (easy)
+(j) DB-P2 (medium)
+(k) DB-P3 (hard)
+(l) DB-P3 (hard)
+Figure 7.
+Comparison of the top five sampling methods for accuracy (a-c, e-g, i-k) and time-accuracy (d, h, l) performance across three node (a-d), three edge (e-h), and three path hypotheses (i-l) for DBLP.
 4.5. Accuracy
-Table 3 presents the accuracy performance of eleven existing samplers and $\text{PHASE}_{\text{opt}}$ at a specific sampling proportion, shown in column three.
-We select this proportion to balance stability and discriminative power, ensuring the accuracy is sufficiently stabilized without saturating across all samplers.
-For each dataset and type of hypothesis, the accuracy is the average accuracy of three hypothesis examples, each derived from an average of 30 runs.
-The highest accuracy in each row is highlighted in bold.
-In Table 3, $\text{PHASE}_{\text{opt}}$ demonstrates robust performance in all cases, except for MovieLens edge hypotheses, in where $\text{PHASE}_{\text{opt}}$ slightly trails behind NBRW.
-Comparing among columns, some random walk based samplers, namely SRW, RWR, and ShortestPathS, exhibit good accuracy performance for edge and path hypothesis.
-On the contrary, RES, RNS, and DBS show ineffectiveness in handling path hypotheses.
-These observations align with their sampling mechanisms.
-These node and edge samplers can hardly preserve the path information from $\mathcal{G}$ in the sampled graph.
-Additionally, comparing $\text{PHASE}_{\text{opt}}$ and FrontierS, $\text{PHASE}_{\text{opt}}$’s superior performance indicates the efficacy of the two weight functions in enhancing the sampler’s hypothesis-awareness to sample relevant nodes, edges, and paths.
-(a) Time $(l=3)$
-(b) Time ($l=4$)
-Figure 6. Time versus accuracy plots for DBLP path hypotheses with length three (a) and four (b).
-(a) DB-N1
-(b) DB-N2
-(c) DB-N3
-(d) DB-N3
-(e) DB-E1
-(f) DB-E2
-(g) DB-E3
-(h) DB-E3
-(i) DB-P1
-(j) DB-P2
-(k) DB-P3
-(l) DB-P3
+Table 3 presents the accuracy of 11 existing samplers and $\text{PHASE}_{\text{opt}}$ with a fix $B$ in column three.
+$B$ is set to ensure the accuracy is sufficiently stabilized without saturating across all samplers.
+Each row reports an average accuracy from three examples, based on 30 runs each.
+The highest and second highest accuracies are highlighted in bold and underlined, respectively.
+Table 3 shows $\text{PHASE}_{\text{opt}}$’s robust performance across most scenarios, except for MovieLens edge hypotheses where it slightly lags behind NBRW.
+SRW, RWR, and ShortestPathS perform well on edge and path hypotheses, whereas RES, RNS, and DBS struggle with path hypotheses.
+This observation aligns with their sampling mechanisms: node and edge samplers can hardly preserve the path information from $\mathcal{G}$.
+Also, $\text{PHASE}_{\text{opt}}$ outperforms FrontierS, showcasing the effectiveness of the two weight functions in enhancing the sampler’s hypothesis-awareness.
+We rank the accuracy of all samplers by averaging their accuracy per column in Table  3 and identify the top five: $\text{PHASE}_{\text{opt}}$, NBRW, ShortestPathS, RWR, and SRW.
+The accuracy performance on individual hypotheses for DBLP and Yelp is shown in Figures  7 and  8, respectively.
+Figures for MovieLens are omitted due to the space constraints and result similarity.
+Subfigures a-c, e-g, and i-k depict accuracy versus $B$ for three node, edge, and path hypotheses, respectively, with truncated x-axes to highlight the convergence.
+(a) YP-N1 (easy)
+(b) YP-N2 (medium)
+(c) YP-N3 (hard)
+(d) YP-N3 (hard)
+(e) YP-E1 (easy)
+(f) YP-E2 (medium)
+(g) YP-E3 (hard)
+(h) YP-E3 (hard)
+(i) YP-P1 (easy)
+(j) YP-P2 (medium)
+(k) YP-P3 (hard)
+(l) YP-P3 (hard)
 Figure 8.
-Comparison of the top five sampling methods for accuracy and time (d, h, l) efficiency across three node (a, b, c), three edge (e, f, g), and three path hypotheses (i, j, k) for the DBLP dataset.
-We establish an accuracy ranking for all samplers by averaging their accuracy performance per column in Table 3 and select the top five: $\text{PHASE}_{\text{opt}}$, NBRW, ShortestPathS, RWR, and SRW.
-These samplers’ accuracy convergence on individual hypotheses is compared in Figures 7 and 8 for DBLP and Yelp datasets, respectively.
-The figure for MovieLens is excluded due to space constraints and the similarity of results.
-In these figures, a-c, e-g, and i-k display the accuracy performance of three node, three edge, and three path hypotheses, respectively.
-Hypotheses of the same type are structured to have progressively fewer relevant nodes, edges, or paths in $\mathcal{G}$.
-The x-axis is truncated to clearly show the convergence.
-In Figures 7 and 8, $\text{PHASE}_{\text{opt}}$ consistently outperforms other methods across all sampling proportions for most hypotheses.
-For hypotheses of the same type, $\text{PHASE}_{\text{opt}}$, with hypothesis-awareness, consistently demonstrates robust accuracy performance for the hypotheses with fewer relevant nodes, edges, or paths in $\mathcal{G}$.
-This is particularly evident for DBLP edge and path hypotheses.
-When relevant nodes, edges, and paths in $\mathcal{G}$ are abundant, conventional sampling methods may have sampled a sufficient number of relevant nodes, edges, or paths to achieve high accuracy, which can diminish the distinctive advantage of $\text{PHASE}_{\text{opt}}$.
-Delving into Figure 7, $\text{PHASE}_{\text{opt}}$ dominates in node and path hypotheses for the DBLP dataset.
-However, ShortestPathS slightly outperforms $\text{PHASE}_{\text{opt}}$ in edge hypotheses one and two, possibly due to the high betweenness centrality of these edges, implying that many shortest paths traverse those edges.
-Moving to Figure 8, $\text{PHASE}_{\text{opt}}$ demonstrates clear superiority across all types of hypotheses for the Yelp dataset.
-Nonetheless, ShortestPathS occasionally exhibits competitive performance compared to $\text{PHASE}_{\text{opt}}$ in path hypotheses, which may also be attributed to the betweenness centrality of those paths in $\mathcal{G}$.
-Figure 5 shows how path length affects accuracy using the DBLP dataset. $\text{PHASE}_{\text{opt}}$ enables faster convergence of accuracy with longer path lengths, outperforming other samplers.
+Comparison of the top five sampling methods for accuracy (a-c, e-g, i-k) and time-accuracy (d, h, l) performance across three node (a-d), three edge (e-h), and three path hypotheses (i-l) for Yelp.
+In Figures 7 and  8, $\text{PHASE}_{\text{opt}}$ consistently outperforms other samplers across most hypotheses and sampling proportions.
+When there are abundant relevant nodes, edges, or paths in $\mathcal{G}$ (e.g, easy and medium hypotheses), hypothesis-agnostic samplers can perform well, reducing $\text{PHASE}_{\text{opt}}$’s relative advantage.
+However, its superiority is more obvious for more difficult hypotheses (e.g., hard hypotheses).
+Also, for difficult hypotheses with longer paths in DBLP, $\text{PHASE}_{\text{opt}}$ achieves the highest accuracy at any sampling proportion, as shown in Figures 6(a) and  6(b).
+Moreover, ShortestPathS sometimes shows competitive performance, possibly due to the high betweenness centrality of the concerned nodes, edges, or paths, which implies many shortest paths traversing them.
 4.6. Scalability
-(a) YP-N1
-(b) YP-N2
-(c) YP-N3
-(d) YP-N3
-(e) YP-E1
-(f) YP-E2
-(g) YP-E3
-(h) YP-E3
-(i) YP-P1
-(j) YP-P2
-(k) YP-P3
-(l) YP-P3
-Figure 8.
-Comparison of the top five sampling methods for accuracy and time (d, h, l) efficiency across three node (a, b, c), three edge (e, f, g), and three path hypotheses (i, j, k) for the Yelp dataset.
-In Table 4, samplers’ execution time increases as the dataset becomes larger.
-Among them, RNS generally exhibits the shortest execution time as it uniformly samples nodes randomly.
-The performance of $\text{PHASE}_{\text{opt}}$ varies across datasets.
-For the Yelp dataset, it ranks among the highest, particularly for the path hypothesis.
-This can be attributed to the limited node types in Yelp, resulting in longer extraction times for condensed relevant paths from $\mathcal{S}$.
-For the DBLP dataset, $\text{PHASE}_{\text{opt}}$ ranks among the lowest five.
-However, instead of considering execution time alone, we should consider both execution time and accuracy performance to evaluate a sampler’s efficiency and effectiveness.
-For the top five samplers ranked by their accuracy performance, we plot their accuracy versus execution time in subfigures d, h, and l of Figures 7 and 8.
-Due to space limitations, only the third hypothesis, which consists of the least number of relevant nodes, edges, or paths in the input graph, of each type is included.
-First of all, we observe that the curve of $\text{PHASE}_{\text{opt}}$ is always at the bottom among others.
-This means that $\text{PHASE}_{\text{opt}}$ consistently achieves high accuracy in the shortest amount of time across all plots.
-Additionally, its execution time does not exhibit exponential growth compared to other existing hypothesis-agnostic samplers.
-This can be attributed to $\text{PHASE}_{\text{opt}}$’s hypothesis-awareness, which allows it to efficiently sample relevant nodes, edges, and paths from the input graph to verify the hypothesis.
-The impact of path length on accuracy and execution time is illustrated in Figure 6.
-As the path length increases, $\text{PHASE}_{\text{opt}}$ consistently achieves the highest accuracy within the shortest execution time.
+In Table  4, the lowest execution time is in bold.
+First, the execution time increases with dataset size.
+Second, RNS generally requires the shortest execution time because it uniformly samples nodes.
+Our method, $\text{PHASE}_{\text{opt}}$, has varying time performance across datasets.
+Specifically, it ranks among the highest in execution time for Yelp, especially for path hypotheses.
+This is attributed to the limited node types that extend path extraction times.
+While for DBLP, its execution time ranks in the lowest five.
+As $\text{PHASE}_{\text{opt}}$ can achieve high accuracy with a small $B$, it is unfair to compare the time efficiency at a fixed $B$.
+Instead, we plot accuracy versus execution time in subfigures d, h, and l of Figures  7 and  8 for DBLP and Yelp, respectively, for the top five samplers ranked by accuracy.
+Due to the space constraints, only the most difficult hypothesis of each type is shown.
+We measure time and accuracy starting from $B=1000$ in increments of 1000 until the accuracy reaches one, or time reaches 50s (DBLP) and 30s (Yelp).
+We find that $\text{PHASE}_{\text{opt}}$ consistently achieves high accuracy in the shortest time and with the least $B$.
+Also, the execution time does not grow exponentially.
+Moreover, when the path length increases, as shown in Figure  6(c) and  6(d), $\text{PHASE}_{\text{opt}}$ achieves the highest accuracy in the shortest time.
 5. Related Work
-There has been a lot of research interest on hypothesis testing (Hunter et al., 2008; Gibbs et al., 2022; Arias-Castro and Verzelen, 2014; Xia and Li, 2017; Croft et al., 2011; Ghoshdastidar and von Luxburg, 2018; Omidvar-Tehrani et al., 2019), proposing sampling techniques on graphs (Leskovec and Faloutsos, 2006; Lee et al., 2012; Li et al., 2019).
-In this section we summarize the most representative works.
+There has been a lot of research interest on hypothesis testing  (Hunter et al., 2008; Gibbs et al., 2022; Arias-Castro and Verzelen, 2014; Xia and Li, 2017; Croft et al., 2011; Ghoshdastidar and von Luxburg, 2018; Omidvar-Tehrani et al., 2019) and graph sampling techniques (Leskovec and Faloutsos, 2006; Lee et al., 2012; Li et al., 2019).
+In this section, we summarize the most representative works.
 Hypothesis Testing on Graphs.
@@
 In (Xia and Li, 2017), nodes are objects of interest, and one-sample hypothesis testing is used to detect conditional dependence between nodes in brain connectivity networks.
-Our goal is more related to works when the object of interest is nodes.
-However, there are two differences between our work and the existing ones.
-First, we enable more expressive hypotheses, including edge and path hypothesis, on graphs.
-Second, we leverage graph sampling methods on graph hypothesis testing.
+Our goal is closely related to works that focus on nodes as the object of interest.
+However, there are two major differences between our work and the existing ones.
+First, we enable more expressive hypotheses, including edge and path hypotheses, on graphs.
+Second, we leverage graph sampling methods to conduct hypothesis testing.
 Sampling Methods on Graphs.
 Most sampling methods are designed to sample a representative subgraph to accurately estimate graph properties (Leskovec and Faloutsos, 2006; Stutzbach et al., 2009; Maiya and Berger-Wolf, 2010; Hübler et al., 2008).
-Leskovec and Faloutsos (Leskovec and Faloutsos, 2006) are the first to study this problem in real-world networks (Leskovec and Faloutsos, 2006) by proposing sampling techniques to maintain degree distribution, clustering coefficient and distribution of component sizes.
+Leskovec and Faloutsos  (Leskovec and Faloutsos, 2006) are the first to study this problem in real-world networks (Leskovec and Faloutsos, 2006) by proposing sampling techniques to maintain degree distribution, clustering coefficient, and distribution of component sizes.
 Hübler et al.
 (Hübler et al., 2008) propose the Metropolis-Hastings sampling method.
-Later, Maiya and Berger-Wolf (Maiya and Berger-Wolf, 2010) propose a community representativeness sample and design a community structure expansion sampler.
-Many sampling methods are proposed to improve the efficiency and convergence rate of simple random walk (Lee et al., 2012; Li et al., 2019).
-Besides representative sampling, other sampling methods are designed for specific tasks, such as graph compression (Besta et al., 2019), community detection(Maiya and Berger-Wolf, 2010), and graph visualization (Rafiei and Curial, 2005; Wu et al., 2017).
+Maiya and Berger-Wolf (Maiya and Berger-Wolf, 2010) define the community representativeness sample and propose a community structure expansion sampler.
+Later, many sampling methods are proposed to improve the efficiency and convergence rate of simple random walk (Lee et al., 2012; Li et al., 2019).
+Besides representative sampling, some sampling methods are designed for specific tasks, such as graph compression (Besta et al., 2019), community detection (Maiya and Berger-Wolf, 2010), and graph visualization (Rafiei and Curial, 2005; Wu et al., 2017).
 However, none of them is designed for hypothesis testing on attributed graphs.
 6. Conclusion
-In this paper, we developed a framework for hypothesis testing on large attributed graphs.
-Our framework accommodates eleven common hypothesis-agnostic samplers as well as new hypothesis-aware samplers.
-We proposed dedicated optimizations to speed up sampling while achieving high test significance and accuracy.
-We demonstrated theoretically and empirically that the convergence rate of our hypothesis-aware sampling.
+In this paper, we develop a framework for hypothesis testing on large attributed graphs, which accommodates 11 existing hypothesis-agnostic samplers and new hypothesis-aware samplers.
+We propose dedicated optimizations to speed up sampling while achieving high test significance and accuracy.
+We also demonstrate theoretically and empirically that our hypothesis-aware sampling achieves earlier convergence of the hypothesis estimator than other methods.
 Our work opens several new directions in the area of hypothesis sampling on graphs.
@@
 The second direction is to handle more expressive hypotheses and specify two-sample and multiple sample scenarios.
+Acknowledgements.
+Reynold Cheng, Yun Wang, and Chrysanthi Kosyfaki were supported by the University of Hong Kong (Project 109000579), the HKU Outstanding Research Student Supervisor Award 2022-23, and the HKU Faculty Exchange Award 2024 (Engineering).